### v1

# Intelligent-Based Neural Networks and Optimal Control of Fractional Order Ebola Virus Dynamics

Abstract
Ebola virus disease is a severe, often deadly form of hemorrhagic fever, and is rapidly transmissible. It’s spread through human contact is increased through the direct exchange of infected blood, body fluids, and organs, as well as contact with surfaces on which these infected fluids have been dropped. Ebola’s complicated dynamics spread beyond integer order models, representing a central intersection for the application of machine learning to control theory. In the present study, we construct an innovative prime framework of fractional order calculus which integrates data-driven machine learning with optimal control in order to model and analyze the Ebola virus disease. We develop an eight-compartment model based on the Caputo fractional derivative to incorporate the non-Markovian memory effects inherent in the disease’s progression and regression. The system explicitly structures the distinct transmission pathways from symptomatic and asymptomatic infections, as well as the critical transmission dynamics associated with deceased individuals. We analytically prove the global well-posedness of the model, compute the basic reproduction number $\mathcal{R}_{0}$ to characterize the epidemic threshold, and prove global stability theorems. Sensitivity analysis indicates that the basic reproduction number is most critically sensitive to transmission rate, incubation period, and the relative infectivity of the deceased. The treatment-safe burial synergy optimization control provides a simultaneous 86.5% of the target morbidity and mortality rate control of the population. The safe burial host intervention is the most effective control. Our study categorizes the best control strategies and orders them in terms of effectiveness. The proposed disease-informed neural network achieves practically perfect predictive accuracy and describe the test set $R^{2}$ values of 0.991 - 0.999, and correspondingly 99.1% - 99.9% prediction accuracy as well as align with the real-world epidemic behavior. The first successful combination of process-based modeling with artificial intelligence for dealing an outbreak of an infectious disease has been established in this work. This work presents an innovative mechanistic modeling framework for infectious diseases, which has the potential to advance epidemic response analysis.
Keywords: Fractional Calculus, Ebola Modeling, Epidemic Threshold ($\mathcal{R}_{0}$), Sensitivity Analysis, Optimal Control, Disease-Informed Neural Networks.
1 Introduction
Ebola Virus Disease (EVD) continues to be one of the most difficult problems for the world health organization to tackle, with estimated case fatality rates varying from concerning 25% to 90% across various outbreaks [1]. The Ebola Virus Disease (EVD) is classified as a tropical zoonosis which involves cross transmission from animal reservoirs to human populations [2, 3]. The earliest recorded case of the disease occurred in Zaire (now known as Democratic Republic of Congo) in 1976, named after the Ebola River where the first outbreak took place [4]. With data around confirmed cases and deaths standing at 28,646 and 11,323, respectively, and spanning through three countries; Liberia, Guinea, and Sierra Leone, the ‘West African Ebola Virus Disease epidemic’ (2014-16) is considered the biggest ‘Ebola’ outbreak in history. The surprising number of cases and deaths during the outbreak and the new ways of fighting the complexity both demonstrated the failure of prediction and outbreak control traditional prediction and control methods used [5]. The outbreak still emphasized the complexity controlled in the actual mathematics and logic of the base prediction and control frameworks prediction, controlling, and modeling control frameworks outlined around the multi and complex dimensions in the spread of the EVD. Mathematical techniques in epidemiology and their application span new prospects and new dimensions in the field [6, 7]. While these techniques were developed in pure science, their fundamental calculations and distributions resource divisions and practical works approximations [8]. The epidemiology of the western African case outbreak hence has triggered the much-needed efficiency flow in prediction case numbers and evaluating the efficiency of preventive measures and their real-time value [9]. The greatest puzzling challenge suffering in the environmental and acoustic scales spread while modalities of EVD transmissions is the combination of factors that various phoneme extents growing paradigm [10]. The epidemiology of Ebola Virus Disease (EVD) has many distinguishing features which imply that estimating the disease outbreaks remaining traditional techniques will not be sufficient. For instance, the disease’s incubation period is highly variable and spans anywhere from 2 to 21 days, with the 8-10 days interval more commonly renowned [11]. There is a further complication within the disease’s clinical presentation and development as initial fever, weakness, and diarrhea unaccompanied by bleeding and other clinical symptoms do not signal an active infection and contagiousness. Additionally, contact with infected animals and the consumption of bush-meat are the first chance for the disease’s human transmission from its natural reservoirs of fruit bats [12, 13]. Subsequently, the disease transmission post- infection is unlinked and random. Furthermore, symptomatic ($I_{s}$), asymptomatic ($I_{a}$), and post-mortem ($D$) transmission routes for the human population. [14]. The high level of infectiousness of dead body, coupled with the perseverance of customarily traditional burial practices in the region, requires mathematically sophisticated treatment beyond typical integer-order differential equations, as outlined in earlier works [15]. Evidence shows that washing the deceased as part of ethnically prescribed performs is a relevant vector for transmission. Besides, the prospect of reinfection with EVD remains very real, as survivors own only quickly protective, low-potency antibodies, with research showing the complete absence of Ebola antibodies in some survivor’s years post-recovery [16, 17]. Such an immunological situation is certainly an additional complication from the modeling perspective whereby the absence of an unified framework for archiving immune waning over a long period rightly remains a gap. The natural, spatial spread of EVD transmission improves so far, another level of complexity to the analytical modelling of the disease dynamics. The dynamics of EVD differ significantly across different geographic scales. Urbanized areas tend to experience outbreaks at a faster rate than rural areas because of the high population density, varied patterns of movement, and better medical infrastructure [18]. Moreover, the phenomenon of superspreader events, where a small proportion of the infected population is responsible for a suspiciously high number of secondary infections, adds a considerable amount of variability and complexity that challenges resolution by purely deterministic approaches [19, 20]. These characteristics need modeling methods that combine large-scale spatial dynamics with fine-scale spatio-temporal transmission patterns [21]. Although traditional modeling techniques such as the SI, SIR, and SEIR frameworks form the basis of understanding, they all fundamentally relate to the Markovian assumptions which do not take into account memory and long-range chronological dependencies which are characteristic of EVD spread [22, 23]. These limitations are particularly clear when attempting to model the distributed incubation periods, the decline of vaccine-induced immunity, and the heterogeneous contact patterns during the real-world epidemic spread [11, 24]. The SEIR model used by [25] and [26], and the SIHDP model employed by [27] and in [28] SEIHFDBR are notable developments, however, they are still limited in the integer-order nature of their models. Fractional calculus has long shown the potential to model memory effects and other non-local connections, and thus provides a framework well-intentioned of study to overcome some of the noted deficiencies: [29]. Unlike integer-order derivatives, which always adopt Markovian processes, fractional operators capture the distributed suspensions which are characteristic of epidemiological systems, such as extended incubation period, waning immunity, and environmental reservoirs of pathogens [30]. Evidence of progress in the modeling of infectious diseases for which data are available and for which fractional calculus applies exhibits greater concord with experimental indication, especially for diseases with long-range chronological correlations and power-law decline in the efficiency of involvements [31, 32]. The non-Markovian nature of some fractional derivatives offers a better-suited mathematical framework for modeling the transmission of disease in and among structured populations [33]. In addressing challenges, some recent studies have included various extensions. For example, [34] included pathways for environmental contamination in their SIRPD model with respect to the Ebola pathogens in the environment. In [35] worked on the decomposition of the deceased compartment by distinguishing dead but not buried, and dead who are buried. Control optimization has been studied by, for instance, [36] on optimal combinations of vaccination and treatment, and [37] on educational campaigns in a two-patch model. Media awareness campaigns have been studied by [38] and [39] and have shown the importance of information broadcasting in controlling a disease. The link of Machine Learning algorithms with Process-based model and models of machine learning with traditional epidemiological forecasting is a new concept as noted by [40]. Predictable models afford a theoretical understanding of epidemiology through biological lenses and as such are weak at capturing sophisticated structures non-linear in nature that are defined with basic compartmental approaches [41]. Recent activities in hybrid frameworks combining the Physics-based model interpretability and the neural networks’ model complexity to capture epidemiological dynamics have demonstrated improved predictive performance [42]. But, as advanced computational techniques apply to EVD, there are gaps such as real time parameter estimation from EVD noise where EVD observation data is incomplete [43]. The development of neural networks with physics-informed neural network (PINN) and epidemiology constraints could bridge this gap [44]. However, modern modeling system contexts still have three main shortcomings. One, they have a limited representation of the memory effects of fractional-order derivatives, which is crucial for modeling the distributed intervals in the progression of a disease and the corresponding delay of the immune response. Two, there is still a lack of fully integrating optimal control theory combined with advanced multi-compartment models that summarize the entirety of the EVD transmission dynamics. Three, here is still insufficient use of machine learning for the estimation of parameters for complicated epidemiological models which reduces the flexibility and accuracy of the models in current practice. The application of optimal control theory is underdeveloped yet central to the design of effective intervention and control strategies for the economic epidemiology systems. Identifying optimal control strategies is a challenging mathematical problem. It requires temporally-resolved control strategies and sophisticated numerical methods for high dimensional, nonlinear systems with memory [45]. Also, minimizing the case count, death, and economic loss for epidemic reduction response requires Pareto-optimal strategies that balance the contradictory objectives [46]. The recent, rapid advancements in adaptable computational optimal control technologies—particularly those dealing with the direct collocation of the control function and the system trajectory and the Hamiltonian-based methods are methods of great attention in computational biologically structured control [47]. The development of fractional calculus has presented new ways to tackle these problems [29]. The Caputo fractional derivative approaches non-local chronological dependencies and memory effects with its non-Markovian structure, which mathematically captures epidemiological data [48]. Preliminary work connecting fractional-order models focused on EVD, such as [49] with the Atangana-Baleanu-Caputo operators and [50] with the nonlinear time-fractional mathematical model. Despite these advances, there is still much to be anticipated in the overlap among these fields, mainly control theory, machine learning, and the other works on the theme. This research discourses these important gaps by revolutionary a new computational framework that combines fractional-order dynamics, multi-pathway transmission architectures, and control optimization, which is enhanced by machine learning. Building upon the preceding work by [34] which did not include exposed compartments, hospitalizations, and safe burial practices, I also consider reinfection through the immunological studies by [16] and [17]. This research directs numerous contributions and novelties:
[…]
  • Comprehensive mathematical analysis signifying the well-posedness of the system through Banach’s fixed-point theorem, the basic reproduction number of the model $\mathcal{R}_{0}$ exogenously resulting through the next generation matrix, and the proof of global stability of the system through fractional Lyapunov methods [48].
  • Derivation and providing an analytical solution for a multi-objective optimal control problem with four intervention strategies (personal-protection, vaccination, treatment, and safe burial) via Pontryagin’s Maximum Principle.
  • Establishing a Disease-Informed Neural Network (DINN) framework for simultaneous solution discovery and parameter estimation which inserts physical constraints for better predictive performances.
  • Detailed cost-effectiveness analysis for optimal intervention portfolios for public health choices.
This model potentially assistances public health authorities in responding to outbreaks more effectively by evidence-based intervention, which can be game-changing in undertaking and enhancing the efficiency of outbreak responses. Integrating mathematical modeling with computational intelligence, this study attempts to productive infectious disease interventions through addressing the organized response and gaps in the current outbreak response system. This paper proceeds as follows. In Section 2, I discuss the elementary ideas of fractional calculus which is key to developing the model. In Section 3, I develop an eight-compartment, fractional-order Ebola transmission model. Section 4 contains the in-depth mathematical theory regarding the model’s well-posedness, equilibrium, and stability over the system. Section 5 conducts global sensitivity analysis of the dominant parameters through Latin Hypercube Sampling and Partial Rank Correlation Coefficients. In Section 6, I formulate and solve the optimal control problem based on Pontryagin’s Maximum Principle. Section 7 incorporates the neural network model concerning disease and the results obtained in the corresponding numerical analysis. In the closing Section 8, I analyze the outcome of the model along with its drawbacks and suggest some directions for further research.
2 Preliminaries
This part sets out the basic principles of fractional calculus which is essential to the construction and analysis of our model.
Definition 2.1 (Caputo Fractional Derivative)
Let $\alpha\in(0,1]$ and $u\in C^{1}([0,T])$. For any function $u$ which is continuously differentiable in the closure $[0,T]$ the Caputo fractional derivative of order $\alpha$ is defined in [29] as follows:
$^{C}D_{t}^{\alpha}u(t)=\frac{1}{\Gamma(1-\alpha)}\int_{0}^{t}(t-\tau)^{-\alpha}u^{\prime}(\tau)d\tau.$
This operator is preferred for modeling dynamical systems as it accommodates standard initial conditions, and $^{C}D_{t}^{\alpha}u(t)\to u^{\prime}(t)$ as $\alpha\to 1^{-}$, recovering the classical integer-order derivative.
Definition 2.2 (Mittag-Leffler Function)
The two-parameter Mittag-Leffler function is defined by the series [29]:
$E_{\alpha,\beta}(z)=\sum_{k=0}^{\infty}\frac{z^{k}}{\Gamma(\alpha k+\beta)},\quad\alpha,\beta>0,\,z\in\mathbb{C}.$
This function generalizes the exponential function and is crucial for constructing solutions to fractional differential equations, playing a role analogous to the exponential in integer-order systems.
Definition 2.3 (Fractional Lyapunov Stability)
The equilibrium point $y^{*}$ of the fractional-order system $^{C}D_{t}^{\alpha}y(t)=f(t,y(t))$ is said to be Mittag-Leffler stable if
$\|y(t)-y^{*}\|\leq\left[m(y(0)-y^{*})E_{\alpha,1}(-\lambda t^{\alpha})\right]^{b},$
where $\alpha\in(0,1)$, $\lambda\geq 0$, $b>0$, $m(0)=0$, and $m(y)\geq 0$. Mittag-Leffler stability implies asymptotic stability [48].
Definition 2.4 (Physics-Informed Neural Network Loss)
For a system governed by $^{C}D_{t}^{\alpha}\mathbf{Y}(t)-\mathbf{F}(t,\mathbf{Y}(t))=0$, the loss function for a Physics-Informed Neural Network (PINN) $NN(t;\theta)$ approximating $\mathbf{Y}(t)$ is defined as:
$\mathcal{L}(\theta)=\mathcal{L}_{data}+\mathcal{L}_{physics},$
where $\mathcal{L}_{data}=\frac{1}{N_{d}}\sum_{i=1}^{N_{d}}|NN(t_{i};\theta)-\mathbf{Y}_{i}|^{2}$ enforces data fidelity, and $\mathcal{L}_{physics}=\frac{1}{N_{c}}\sum_{j=1}^{N_{c}}|{}^{C}D_{t}^{\alpha}NN(t_{j};\theta)-\mathbf{F}(t_{j},NN(t_{j};\theta))|^{2}$ enforces the physical law described by the fractional differential equation.
Theorem 2.1 (Stability Condition for Fractional Systems)
Let $\alpha\in(0,1]$. An equilibrium point $y^{*}$ of the autonomous system $^{C}D_{t}^{\alpha}y(t)=f(y(t))$ is asymptotically stable if all eigenvalues $\lambda_{i}$ of the Jacobian matrix $J=\partial f/\partial y|_{y^{*}}$ satisfy the condition [48]:
$|\arg(\lambda_{i})|>\frac{\alpha\pi}{2},\quad\forall i.$
This condition generalizes the classical Routh-Hurwitz criterion to fractional-order systems.
3 Fractional-Order Ebola Transmission Dynamics
We formulate a Caputo fractional-order model ($\alpha\in(0,1]$) which incorporates the multi-pathway transmission, memory effects, and time-varying interventions of Ebola. The governing system is given as follows.
${}^{C}D_{t}^{\alpha}S(t)$ $=\Lambda-\frac{c\beta S(I_{s}+\eta_{a}I_{a}+\eta_{d}D)}{N}+\omega V-(v+\mu)S,$ (1)
${}^{C}D_{t}^{\alpha}V(t)$ $=vS-(1-\varepsilon)\frac{c\beta V(I_{s}+\eta_{a}I_{a}+\eta_{d}D)}{N}-(\mu+\omega)V,$
${}^{C}D_{t}^{\alpha}E(t)$ $=\frac{c\beta[S+(1-\varepsilon)V](I_{s}+\eta_{a}I_{a}+\eta_{d}D)}{N}-(\mu+\sigma)E,$
${}^{C}D_{t}^{\alpha}I_{s}(t)$ $=p\sigma E-(\gamma_{s}+\delta_{s}+h_{s}+\mu)I_{s},$
${}^{C}D_{t}^{\alpha}I_{a}(t)$ $=(1-p)\sigma E-(\gamma_{a}+\mu)I_{a},$
${}^{C}D_{t}^{\alpha}H(t)$ $=h_{s}I_{s}-(\gamma_{h}+\delta_{h}+\mu)H,$
${}^{C}D_{t}^{\alpha}D(t)$ $=\delta_{s}I_{s}+\delta_{h}H-\mu_{d}D,$
${}^{C}D_{t}^{\alpha}R(t)$ $=\gamma_{s}I_{s}+\gamma_{a}I_{a}+\gamma_{h}H-\mu R.$
The force of infection is given by:
$\lambda=\frac{c\beta(I_{s}+\eta_{a}I_{a}+\eta_{d}D)}{N},$
and for notational simplicity, we define:
$q_{0}$ $=(v+\mu),$ $q_{1}$ $=(\mu+\omega),$ $q_{2}$ $=(\mu+\sigma),$
$q_{3}$ $=(\gamma_{s}+\delta_{s}+h_{s}+\mu),$ $q_{4}$ $=(\gamma_{a}+\mu),$ $q_{5}$ $=(\gamma_{h}+\delta_{h}+\mu),$
$q_{6}$ $=\mu_{d}.$
The model parameters include: natural mortality ($\mu$) and recruitment ($\Lambda$); vaccination rate $v(t)$ with waning immunity ($\omega$); and force of infection $\lambda(t)$ driving exposures. Exposed individuals ($E$) progress to infectious states at rate $\sigma$, splitting into symptomatic ($I_{s}$, proportion $p$) and asymptomatic ($I_{a}$, 1p1-p) pathways. Symptomatic cases face hospitalization ($h_{s}$), disease-induced mortality ($\delta_{s}$), or recovery ($\gamma_{s}$). Asymptomatic and hospitalized individuals recover at rates $\gamma_{a}$ and $\gamma_{h}$ respectively, with hospital mortality ($\delta_{h}$). Deceased individuals ($D$) transmit infection until safe burial ($\mu_{d}$). Transmission coefficients $\beta$, $\eta_{a}$, and $\eta_{d}$ govern infectivity from symptomatic, asymptomatic, and post-mortem sources. The system is governed by:
${}^{C}D_{t}^{\alpha}S(t)$ $=\Lambda-\lambda S+\omega V-q_{0}S,$ (2)
${}^{C}D_{t}^{\alpha}V(t)$ $=vS-(1-\varepsilon)\lambda V-q_{1}V,$
${}^{C}D_{t}^{\alpha}E(t)$ $=\lambda[S+(1-\varepsilon)V]-q_{2}E,$
${}^{C}D_{t}^{\alpha}I_{s}(t)$ $=p\sigma E-q_{3}I_{s},$
${}^{C}D_{t}^{\alpha}I_{a}(t)$ $=(1-p)\sigma E-q_{4}I_{a},$
${}^{C}D_{t}^{\alpha}H(t)$ $=h_{s}I_{s}-q_{5}H,$
${}^{C}D_{t}^{\alpha}D(t)$ $=\delta_{s}I_{s}+\delta_{h}H-\mu_{d}D,$
${}^{C}D_{t}^{\alpha}R(t)$ $=\gamma_{s}I_{s}+\gamma_{a}I_{a}+\gamma_{h}H-\mu R.$
The fractional-order system (2) is analyzed with the following initial conditions:
$S(0)=S_{0}\geq 0,\quad V(0)=V_{0}\geq 0,\quad E(0)=E_{0}\geq 0,\quad I_{s}(0)=I_{s_{0}}\geq 0,$
$I_{a}(0)=I_{a_{0}}\geq 0,\quad H(0)=H_{0}\geq 0,\quad D(0)=D_{0}\geq 0,\quad R(0)=R_{0}\geq 0.$
In this Caputo fractional formulation, the system in (2) captures the dynamics of Ebola transmission in an eight-dimensional frame. This construction involves ‘epidemic’ components such as: Susceptible ($S$), Vaccinated ($V$), Exposed ($E$), Symptomatically ($I_{s}$) and Asymptomatically ($I_{a}$) infectious, Hospitalized ($H$), deceased ($D$), and Recovery ($R$) stages, thus integrating critical structure of underlying dynamics. Furthermore, the model participates the fractional-order system of differential equations, thus inserting memory effects.
[…]
Figure 1: Transmission architecture of the fractional Ebola model.
Table 1: Epidemiological state variables and parameter estimates for the fractional-order Ebola model.
Variable Epidemiological Definition Estimated Value
$S(t)$ Susceptible individuals
$V(t)$ Vaccinated individuals
$E(t)$ Exposed (infected, not infectious)
$I_{s}(t)$ Symptomatic infectious individuals
$I_{a}(t)$ Asymptomatic infectious individuals
$H(t)$ Hospitalized/isolated individuals
$D(t)$ Deceased (infectious pre-burial)
$R(t)$ Recovered individuals
$N(t)$ Total living population
$\beta$ Transmission rate 0.287
$\eta_{a}$ Relative infectiousness (asymptomatic) 0.523
$\eta_{d}$ Relative infectiousness (deceased) 0.734
$\sigma$ Incubation rate (day-1) 0.094
$p$ Proportion symptomatic 0.712
$\gamma_{s}$ Recovery rate (symptomatic) 0.068
$\delta_{s}$ Death rate (symptomatic) 0.103
$h_{s}$ Hospitalization rate 0.312
$\gamma_{a}$ Recovery rate (asymptomatic) 0.089
$\delta_{h}$ Death rate (hospitalized) 0.067
Table 2: Parameter ranges and biological interpretation for the fractional-order Ebola model.
Symbol Biological Interpretation Range Source
$\Lambda$ Recruitment rate 100-1000 [1]
$\mu$ Natural mortality rate $3.5\times 10^{-5}$ [51]
$c(t)$ Contact rate function 0.4-1.2 [9]
$v(t)$ Vaccination rate 0.005-0.08 [52]
$\varepsilon$ Vaccine efficacy 0.85-0.95 [52]
$\omega$ Waning immunity rate 1/365-1/270 [53]
$\mu_{d}$ Burial rate 0.15-0.35 [54]
$\alpha$ Fractional order 0.75-0.95 [29]
Figure 2: Model validation: Curve fitting to confirmed Ebola cases ($R^{2}=0.96$, RMSE = 23.45).
4 Mathematical Analysis
4.1 Well-Posedness of the Model
We establish the mathematical well-posedness of the fractional-order Ebola system through existence and uniqueness analysis.
Theorem 4.1 (Existence and Uniqueness)
Given any $\alpha\in(0,1]$, and initial conditions
$\mathbf{X}_{0}=(S_{0},V_{0},E_{0},I_{s0},I_{a0},H_{0},D_{0},R_{0})\in\mathbb{R}_{+}^{8}$, the fractional-order system (2) has the unique solution $\mathbf{X}(t)=(S(t),V(t),E(t),I_{s}(t),I_{a}(t),H(t),D(t),R(t))$ for all $t\geq 0$.
Proof.  Define the state vector $\mathbf{X}(t)=(S,V,E,I_{s},I_{a},H,D,R)^{T}$ and express the system as:
$^{C}D_{t}^{\alpha}\mathbf{X}(t)=\mathbf{F}(t,\mathbf{X}(t)),\quad\mathbf{X}(0)=\mathbf{X}_{0}.$
The equivalent Volterra integral formulation is:
$\mathbf{X}(t)=\mathbf{X}_{0}+\frac{1}{\Gamma(\alpha)}\int_{0}^{t}(t-s)^{\alpha-1}\mathbf{F}(s,\mathbf{X}(s))ds.$ (3)
Assume the operator $\mathcal{T}:C([0,T],\mathbb{R}^{8})\to C([0,T],\mathbb{R}^{8})$ defined by:
$(\mathcal{T}\mathbf{X})(t)=\mathbf{X}_{0}+\frac{1}{\Gamma(\alpha)}\int_{0}^{t}(t-s)^{\alpha-1}\mathbf{F}(s,\mathbf{X}(s))ds.$
For $\mathbf{X},\mathbf{Y}\in C([0,T],\mathbb{R}^{8})$ with $\|\mathbf{X}\|=\sup_{t\in[0,T]}\|\mathbf{X}(t)\|_{\infty}$, the Lipschitz continuity of $\mathbf{F}$ follows from bounded parameters and finite state variables over compact intervals. Specifically, there exists $L>0$ such that:
$\|\mathbf{F}(t,\mathbf{X})-\mathbf{F}(t,\mathbf{Y})\|_{\infty}\leq L\|\mathbf{X}-\mathbf{Y}\|_{\infty}.$
The contraction estimate yields:
$\|\mathcal{T}\mathbf{X}-\mathcal{T}\mathbf{Y}\|\leq\frac{L\tau^{\alpha}}{\Gamma(\alpha+1)}\|\mathbf{X}-\mathbf{Y}\|.$
Selecting $\tau>0$ such that $\frac{L\tau^{\alpha}}{\Gamma(\alpha+1)}<1$ ensures $\mathcal{T}$ is a contraction on $C([0,\tau],\mathbb{R}^{8})$. Banach’s fixed point theorem guarantees a unique local solution $\mathbf{X}^{*}\in C([0,\tau],\mathbb{R}^{8})$.
Global extension to $0,\infty)$ follows from the linear growth condition $\|\mathbf{F}(t,\mathbf{X})\|_{\infty}\leq M(1+\|\mathbf{X}\|_{\infty})$ for some $M>0$, which precludes finite-time blow-up. Iterative continuation establishes global existence and uniqueness. $\square$
4.2 Invariant Region and Attractivity
The fractional-order Ebola system (2) evolves within the biologically feasible region:
$\Omega=\left\{(S,V,E,I_{s},I_{a},H,D,R)\in\mathbb{R}^{8}_{\geq 0}:\quad D\leq\frac{(\delta_{s}+\delta_{h})}{\mu_{d}}\left(\frac{\Lambda}{\mu}+\epsilon\right),N_{L}\leq(\frac{\Lambda}{\mu}+\epsilon)\right\},$ (4)
for $\epsilon>0$, where $N_{L}=S+V+E+I_{s}+I_{a}+H+R$ denotes the living population.
Theorem 4.2 (Invariance and Attractivity)
The region $\Omega$ is:
  1. (i)
    Positively invariant under (2),
  2. (ii)
    Globally attractive in $\mathbb{R}^{8}_{\geq 0}$.
Proof.  (i) Positivity and Invariance: Applying the fractional positivity lemma [29], we verify boundary behavior:
${}^{C}D_{t}^{\alpha}S|_{S=0}=\Lambda+\omega V\geq 0,\quad{}^{C}D_{t}^{\alpha}V|_{V=0}=vS\geq 0,$
${}^{C}D_{t}^{\alpha}E|_{E=0}=\frac{c\beta[S+(1-\varepsilon)V](I_{s}+\eta_{a}I_{a}+\eta_{d}D)}{N}\geq 0,$
${}^{C}D_{t}^{\alpha}I_{s}|_{I_{s}=0}=p\sigma E\geq 0,\quad{}^{C}D_{t}^{\alpha}I_{a}|_{I_{a}=0}=(1-p)\sigma E\geq 0,$
${}^{C}D_{t}^{\alpha}H|_{H=0}=h_{s}I_{s}\geq 0,\quad{}^{C}D_{t}^{\alpha}D|_{D=0}=\delta_{s}I_{s}+\delta_{h}H\geq 0,$
${}^{C}D_{t}^{\alpha}R|_{R=0}=\gamma_{s}I_{s}+\gamma_{a}I_{a}+\gamma_{h}H\geq 0.$
Non-negativity of all boundary fluxes ensures $\mathbb{R}_{+}^{8}$ is invariant. (ii) Boundedness and Attractivity: For the living population $N_{L}$:
$^{C}D_{t}^{\alpha}N_{L}=\Lambda-\mu N_{L}-(\delta_{s}I_{s}+\delta_{h}H)\leq\Lambda-\mu N_{L}.$
Laplace transformation yields:
$N_{L}(s)\leq\frac{\Lambda}{s(s^{\alpha}+\mu)}+N_{L}(0)\frac{s^{\alpha-1}}{s^{\alpha}+\mu},$
with inverse transform:
$N_{L}(t)\leq N_{L}(0)E_{\alpha,1}(-\mu t^{\alpha})+\Lambda t^{\alpha}E_{\alpha,\alpha+1}(-\mu t^{\alpha}).$
As $t\to\infty$, $E_{\alpha,1}(-\mu t^{\alpha})\to 0$ and $t^{\alpha}E_{\alpha,\alpha+1}(-\mu t^{\alpha})\to 1/\mu$, giving:
$\limsup_{t\to\infty}N_{L}(t)\leq\frac{\Lambda}{\mu}.$
For deceased individuals:
$^{C}D_{t}^{\alpha}D\leq(\delta_{s}+\delta_{h})\left(\frac{\Lambda}{\mu}+\epsilon\right)-\mu_{d}D,$
implying:
$\limsup_{t\to\infty}D(t)\leq\frac{\delta_{s}+\delta_{h}}{\mu_{d}}\left(\frac{\Lambda}{\mu}+\epsilon\right).$
Thus, all trajectories are eventually absorbed into $\Omega$. $\square$
4.3 EVD-Free Equilibrium
The state wherein a population exists without any infections of the Ebola virus is referred to the Ebola-free equilibrium (DFE), at which point all infected compartments have completely dissipated:
$E=I_{s}=I_{a}=H=D=0.$
The system (2) reduces to:
$0$ $=\Lambda-q_{0}S+\omega V,$ (5)
$0$ $=vS-q_{1}V,$ (6)
$0$ $=-\mu R.$ (7)
From (7), $R^{*}=0$. Equations (5) and (6) form the linear system:
$q_{0}S-\omega V$ $=\Lambda,$ (8)
$-vS+q_{1}V$ $=0.$ (9)
The coefficient matrix
$A=\begin{pmatrix}q_{0}&-\omega\\ -v&q_{1}\end{pmatrix}$
has determinant
$\det(A)=q_{0}q_{1}-\omega v=\mu(q_{1}+v)>0,$
ensuring a unique solution. Applying Cramer’s rule:
$S^{*}=\frac{\Lambda q_{1}}{\mu(q_{1}+v)},\quad V^{*}=\frac{\Lambda v}{\mu(q_{1}+v)}.$
Thus, the disease-free equilibrium is:
$\mathcal{E}_{0}=\left(\frac{\Lambda q_{1}}{\mu(q_{1}+v)},\frac{\Lambda v}{\mu(q_{1}+v)},0,0,0,0,0,0\right).$
This biologically consistent DFE provides the foundation for stability analysis and reproduction number derivation.
[…]
$\mathcal{R}_{0}$ is computed via the next generation matrix method as expressed in [55]. In the case of the infected subsystem $X=(E,I_{s},I_{a},H,D)^{T}$, the new infection ($F$) and transition ($V$) matrices are as follows:
$F=\begin{pmatrix}0&\beta\Psi&\beta\eta_{a}\Psi&0&\beta\eta_{d}\Psi\\ 0&0&0&0&0\\ 0&0&0&0&0\\ 0&0&0&0&0\\ 0&0&0&0&0\end{pmatrix},\quad V=\begin{pmatrix}q_{2}&0&0&0&0\\ -p\sigma&q_{3}&0&0&0\\ -(1-p)\sigma&0&q_{4}&0&0\\ 0&-h_{s}&0&q_{5}&0\\ 0&-\delta_{s}&0&-\delta_{h}&q_{6}\end{pmatrix}$
where $\Psi=\frac{S^{*}+(1-\varepsilon)V^{*}}{N^{*}}$ with $N^{*}=\frac{\Lambda}{\mu}$.
The inverse transition matrix is:
$V^{-1}=\begin{pmatrix}\frac{1}{q_{2}}&0&0&0&0\\ \frac{p\sigma}{q_{2}q_{3}}&\frac{1}{q_{3}}&0&0&0\\ \frac{(1-p)\sigma}{q_{2}q_{4}}&0&\frac{1}{q_{4}}&0&0\\ \frac{p\sigma h_{s}}{q_{2}q_{3}q_{5}}&\frac{h_{s}}{q_{3}q_{5}}&0&\frac{1}{q_{5}}&0\\ \frac{p\sigma(\delta_{s}q_{5}+\delta_{h}h_{s})}{q_{2}q_{3}q_{5}q_{6}}&\frac{\delta_{s}q_{5}+\delta_{h}h_{s}}{q_{3}q_{5}q_{6}}&0&\frac{\delta_{h}}{q_{5}q_{6}}&\frac{1}{q_{6}}\end{pmatrix}$
The spectral radius of the next generation matrix is $K=FV^{-1}$ yields:
$\mathcal{R}_{0}=\beta\Psi\left[\frac{p\sigma}{q_{2}}\left(\frac{1}{q_{3}}+\frac{\eta_{d}(\delta_{s}q_{5}+\delta_{h}h_{s})}{q_{3}q_{5}q_{6}}\right)+\frac{\eta_{a}(1-p)\sigma}{q_{2}q_{4}}\right]$
Substituting $\Psi$ provides the explicit form:
$\mathcal{R}_{0}=\frac{\beta\sigma[S^{*}+(1-\varepsilon)V^{*}]}{(\Lambda/\mu)q_{2}}\left[\frac{p}{q_{3}}\left(1+\frac{\eta_{d}(\delta_{s}q_{5}+\delta_{h}h_{s})}{q_{5}q_{6}}\right)+\frac{\eta_{a}(1-p)}{q_{4}}\right]$
This threshold quantity governs disease invasion and extinction dynamics.
(a) Cumulative confirmed cases progression
(b) Cumulative deaths progression
(c) Active cases progression
(d) Comparative analysis of all metrics
Figure 3: Ebola Virus Disease outbreak analysis during the 2014-2016 West Africa outbreak showing: (a) cumulative confirmed cases, (b) cumulative deaths, (c) active cases, and (d) comparative analysis of all outbreak metrics.
4.5 Critical Vaccination Coverage
The critical vaccination coverage defines the minimum population immunity required for disease elimination, occurring when $\mathcal{R}_{0}=1$. Beginning with the reproduction number:
$\mathcal{R}_{0}(v)=\mathcal{R}_{0}(0)\cdot\frac{q_{1}+(1-\varepsilon)v}{q_{1}+v}$
where $\mathcal{R}_{0}(0)=K$ represents the vaccination-free reproduction number. Setting $\mathcal{R}_{0}(v_{c})=1$ yields:
$1=\mathcal{R}_{0}(0)\cdot\frac{q_{1}+(1-\varepsilon)v_{c}}{q_{1}+v_{c}}$
Solving for the critical vaccination rate $v_{c}$:
$q_{1}+v_{c}$ $=\mathcal{R}_{0}(0)[q_{1}+(1-\varepsilon)v_{c}]$
$v_{c}-\mathcal{R}_{0}(0)(1-\varepsilon)v_{c}$ $=\mathcal{R}_{0}(0)q_{1}-q_{1}$
$v_{c}[1-\mathcal{R}_{0}(0)(1-\varepsilon)]$ $=q_{1}[\mathcal{R}_{0}(0)-1]$
$v_{c}=\frac{q_{1}[\mathcal{R}_{0}(0)-1]}{1-(1-\varepsilon)\mathcal{R}_{0}(0)}$
The critical vaccination coverage $p_{c}$, representing the proportion vaccinated at equilibrium, is:
$p_{c}=\frac{v_{c}}{q_{1}+v_{c}}=\frac{\mathcal{R}_{0}(0)-1}{\varepsilon\mathcal{R}_{0}(0)}$
Thus, the herd immunity threshold for Ebola elimination is:
$p_{c}=\max\left(0,1-\frac{1}{\varepsilon\mathcal{R}_{0}(0)}\right)$
This establishes the minimal vaccination coverage required to interrupt disease transmission.
Figure 4: 3D surface: Vaccination and waning immunity impact on $\mathcal{R}_{0}$. The surface shows how the basic reproduction number varies with vaccination rate ($v$) and waning immunity rate ($\omega$).
4.6 Local Stability Analysis of the EVD-Free Equilibrium
Theorem 4.3
When $\mathcal{R}_{0}<1$, the EVD-free equilibrium $\mathcal{E}_{0}$ of the system (2) is achieved and remains locally stable; however, it becomes unsteady and $\mathcal{R}_{0}>1$ will prove it to be unstable.
Proof.  The Jacobian matrix evaluated at $\mathcal{E}_{0}$ takes the block triangular form:
$J(\mathcal{E}_{0})=\begin{pmatrix}J_{11}&J_{12}\\ 0&J_{22}\end{pmatrix}$
where $J_{11}$ corresponds to the uninfected compartments $(S,V,R)$ and $J_{22}$ to the infected compartments $(E,I_{s},I_{a},H,D)$. The uninfected subsystem matrix is:
$J_{11}=\begin{pmatrix}-(v+\mu)&\omega&0\\ v&-(\mu+\omega)&0\\ 0&0&-\mu\end{pmatrix}$
with eigenvalues $\lambda_{1}=-\mu$, $\lambda_{2}=-\mu$, $\lambda_{3}=-(\mu+v+\omega)$. All eigenvalues satisfy $|\arg(\lambda_{i})|=\pi>\frac{\alpha\pi}{2}$ for $\alpha\in(0,1]$, confirming stability of the uninfected subsystem. The infected subsystem Jacobian is:
$J_{22}=\begin{pmatrix}-q_{2}&\beta\Psi&\beta\eta_{a}\Psi&0&\beta\eta_{d}\Psi\\ p\sigma&-q_{3}&0&0&0\\ (1-p)\sigma&0&-q_{4}&0&0\\ 0&h_{s}&0&-q_{5}&0\\ 0&\delta_{s}&0&\delta_{h}&-q_{6}\end{pmatrix}$
where $\Psi=\frac{S^{*}+(1-\varepsilon)V^{*}}{N^{*}}$. We decompose $J_{22}=F-V$, where $F$ and $V$ are the next-generation matrices previously defined. The stability of $J_{22}$ is determined by the spectral radius $\rho(FV^{-1})=\mathcal{R}_{0}$. According to the fractional Routh-Hurwitz criteria and properties of Metzler matrices, when $\mathcal{R}_{0}<1$, all eigenvalues of $J_{22}$ have negative real parts and satisfy the fractional stability condition $|\arg(\lambda)|>\frac{\alpha\pi}{2}$.
Conversely, when $\mathcal{R}_{0}>1$, the spectral abscissa of $J_{22}$ is positive, indicating at least one eigenvalue with positive real part. By the fractional stability theorem, this establishes instability of the disease-free equilibrium. $\square$
4.7 Global Stability of EVD-Free Equilibrium
Theorem 4.4
When $\mathcal{R}_{0}<1$, the EVD-free equilibrium $\mathcal{E}_{0}$ of the system (2) is globally asymptotically stable in $\Omega$.
Proof.  Consider the function for the infected compartments already introduced in the above section.
$L(E,I_{s},I_{a},H,D)=E+AI_{s}+BI_{a}+CH+FD,$
where the parameters A, B, C, and F are positive and not arbitrary. To investigate the function, we follow the trajectories described in the system and compute the Caputo derivative along each trajectory:
${}^{C}D_{t}^{\alpha}L=$ $\left[[\frac{c\beta[S+(1-\varepsilon)V](I_{s}+\eta_{a}I_{a}+\eta_{d}D)}{N}]-q_{2}E\right]$
$+A\left(p\sigma E-q_{3}I_{s}\right)+B((1-p)\sigma E-q_{4}I_{a})$
$+C\left(h_{s}I_{s}-q_{5}H\right)+F(\delta_{s}I_{s}+\delta_{h}H-q_{6}D)$
Rearranging terms by compartment:
${}^{C}D_{t}^{\alpha}L=$ $\left[-q_{2}+Ap\sigma+B(1-p)\sigma\right]E$
$+\left[\frac{c\beta[S+(1-\varepsilon)V]}{N}-Aq_{3}+Ch_{s}+F\delta_{s}\right]I_{s}$
$+\left[\frac{c\beta\eta_{a}[S+(1-\varepsilon)V]}{N}-Bq_{4}\right]I_{a}$
$+\left[-Cq_{5}+F\delta_{h}\right]H$
$+\left[\frac{c\beta\eta_{d}[S+(1-\varepsilon)V]}{N}-Fq_{6}\right]D$
We select coefficients to eliminate positive terms:
$C$ $=\frac{F\delta_{h}}{q_{5}},\quad B=\frac{\beta\eta_{a}\Psi}{q_{4}},\quad F=\frac{\beta\eta_{d}\Psi}{q_{6}}$
$A$ $=\frac{\beta\Psi}{q_{3}}+\frac{Ch_{s}}{q_{3}}+\frac{F\delta_{s}}{q_{3}}=\frac{\beta\Psi}{q_{3}}+\frac{F\delta_{h}h_{s}}{q_{3}q_{5}}+\frac{F\delta_{s}}{q_{3}}$
Using the bound $\frac{S+(1-\varepsilon)V}{N}\leq\Psi$ valid throughout $\Omega$, we obtain:
${}^{C}D_{t}^{\alpha}L\leq$ $\left[-q_{2}+\sigma\left(\frac{p\beta\Psi}{q_{3}}+\frac{pF\delta_{h}h_{s}}{q_{3}q_{5}}+\frac{pF\delta_{s}}{q_{3}}+\frac{(1-p)\beta\eta_{a}\Psi}{q_{4}}\right)\right]E$
$+\beta\Psi\left[1-1\right]I_{s}+\beta\eta_{a}\Psi\left[1-1\right]I_{a}+\beta\eta_{d}\Psi\left[1-1\right]D$
Substituting $F=\frac{\beta\eta_{d}\Psi}{q_{6}}$ yields:
$^{C}D_{t}^{\alpha}L\leq q_{2}\left[-1+\mathcal{R}_{0}\right]E$
For $\mathcal{R}_{0}\leq 1$, we have $^{C}D_{t}^{\alpha}L\leq 0$ throughout $\Omega$, with equality if and only if $E=0$. When $E=0$, the system dynamics ensure:
${}^{C}D_{t}^{\alpha}I_{s}$ $=-q_{3}I_{s}\Rightarrow I_{s}\to 0$
${}^{C}D_{t}^{\alpha}I_{a}$ $=-q_{4}I_{a}\Rightarrow I_{a}\to 0$
${}^{C}D_{t}^{\alpha}H$ $=h_{s}I_{s}-q_{5}H\Rightarrow H\to 0$
${}^{C}D_{t}^{\alpha}D$ $=\delta_{s}I_{s}+\delta_{h}H-q_{6}D\Rightarrow D\to 0$
Thus, the maximal invariant set where $^{C}D_{t}^{\alpha}L=0$ is precisely $\mathcal{E}_{0}$. Application of the fractional LaSalle invariance principle establishes global asymptotic stability when $\mathcal{R}_{0}\leq 1$. For $\mathcal{R}_{0}>1$, instability follows from local analysis, while uniform persistence can be demonstrated via established methods in fractional dynamical systems [48]. $\square$
4.8 Endemic Equilibrium Analysis
Theorem 4.5 (Existence of Endemic Equilibrium)
The fractional-order Ebola system (2) admits a unique endemic equilibrium point $\mathcal{E}^{*}=(S^{*},V^{*},E^{*},I_{s}^{*},I_{a}^{*},H^{*},D^{*},R^{*})$ with all components strictly positive if and only if $\mathcal{R}_{0}>1$.
Proof.  At equilibrium, the time derivatives vanish. The force of infection is:
$\lambda^{*}=\frac{c\beta(I_{s}^{*}+\eta_{a}I_{a}^{*}+\eta_{d}D^{*})}{N^{*}}$
The equilibrium equations provide:
$0$ $=\Lambda-\lambda^{*}S^{*}+\omega V^{*}-q_{0}S^{*}$ (10)
$0$ $=vS^{*}-(1-\varepsilon)\lambda^{*}V^{*}-q_{1}V^{*}$ (11)
$0$ $=\lambda^{*}[S^{*}+(1-\varepsilon)V^{*}]-q_{2}E^{*}$ (12)
$0$ $=p\sigma E^{*}-q_{3}I_{s}^{*}$ (13)
$0$ $=(1-p)\sigma E^{*}-q_{4}I_{a}^{*}$ (14)
$0$ $=h_{s}I_{s}^{*}-q_{5}H^{*}$ (15)
$0$ $=\delta_{s}I_{s}^{*}+\delta_{h}H^{*}-q_{6}D^{*}$ (16)
$0$ $=\gamma_{s}I_{s}^{*}+\gamma_{a}I_{a}^{*}+\gamma_{h}H^{*}-\mu R^{*}$ (17)
Expressing infected states in terms of $E^{*}$ as:
$I_{s}^{*}$ $=\frac{p\sigma}{q_{3}}E^{*}\ and\quad I_{a}^{*}=\frac{(1-p)\sigma}{q_{4}}E^{*}$
$H^{*}$ $=\frac{p\sigma h_{s}}{q_{3}q_{5}}E^{*}and\quad D^{*}=\frac{p\sigma}{q_{3}q_{6}}\left(\delta_{s}+\frac{\delta_{h}h_{s}}{q_{5}}\right)E^{*}$
The force of infection becomes:
$\lambda^{*}=\frac{c\beta\sigma E^{*}}{N^{*}}\left[\frac{p}{q_{3}}+\frac{\eta_{a}(1-p)}{q_{4}}+\frac{\eta_{d}p}{q_{3}q_{6}}\left(\delta_{s}+\frac{\delta_{h}h_{s}}{q_{5}}\right)\right]$
Define the composite parameter:
$M=c\beta\sigma\left[\frac{p}{q_{3}}+\frac{\eta_{a}(1-p)}{q_{4}}+\frac{\eta_{d}p}{q_{3}q_{6}}\left(\delta_{s}+\frac{\delta_{h}h_{s}}{q_{5}}\right)\right]$
yielding $\lambda^{*}=ME^{*}/N^{*}$. From the exposed compartment equilibrium:
$E^{*}=\frac{\lambda^{*}[S^{*}+(1-\varepsilon)V^{*}]}{q_{2}},$
Uniting these relations gives the key equation:
$1=\frac{M}{q_{2}N^{*}}[S^{*}+(1-\varepsilon)V^{*}]$
Solving the susceptible-vaccinated subsystem:
$S^{*}$ $=\frac{\Lambda(q_{1}+(1-\varepsilon)\lambda^{*})}{\Delta(\lambda^{*})},\quad V^{*}=\frac{\Lambda v}{\Delta(\lambda^{*})}$
where $\Delta(\lambda^{*})=\left[(\omega+\mu+(1-\varepsilon)\lambda^{*})(\mu+v+\lambda^{*})\right]-\omega v$. This leads to $N^{*}=\Lambda/\mu$ hence the key equation becomes:
$1=\left[\frac{M}{q_{2}(\Lambda/\mu)}\right][S^{*}+(1-\varepsilon)V^{*}]$
At the disease-free state ($\lambda^{*}=0$), this expression equals $\mathcal{R}_{0}$. The monotonic decrease of $S^{*}+(1-\varepsilon)V^{*}]$ with $\lambda^{*}$ guarantees a unique positive solution $\lambda^{*}>0$ when $\mathcal{R}_{0}>1$, establishing the endemic equilibrium. $\square$
4.9 Local Stability Analysis of the Endemic Equilibrium
Theorem 4.6
Consider system (2). When $\mathcal{R}_{0}>1$, the endemic equilibrium $\mathcal{E}^{*}$ is locally asymptotically stable.
Proof.  The Jacobian matrix evaluated at $\mathcal{E}^{*}$ exhibits the block structure:
$J(\mathcal{E}^{*})=\begin{pmatrix}J_{11}&J_{12}\\ J_{21}&J_{22}\end{pmatrix}$
with components:
$J_{11}$ $=\begin{pmatrix}-(\lambda^{*}+q_{0})&\omega\\ v&-\lambda^{*}+\varepsilon\lambda^{*}-q_{1})\end{pmatrix}$,
$J_{12}$ $=\begin{pmatrix}0&-\beta S^{*}\Phi&-\beta\eta_{a}S^{*}\Phi&0&-\beta\eta_{d}S^{*}\Phi&0\\ 0&-\beta(1-\varepsilon)V^{*}\Phi&-\beta\eta_{a}(1-\varepsilon)V^{*}\Phi&0&-\beta\eta_{d}(1-\varepsilon)V^{*}\Phi&0\end{pmatrix}$,
$J_{21}$ $=\begin{pmatrix}\lambda^{*}&(1-\varepsilon)\lambda^{*}\\ 0&0\\ 0&0\\ 0&0\\ 0&0\\ 0&0\end{pmatrix}$,
$J_{22}$ $=\begin{pmatrix}-q_{2}&\beta\Psi\Phi&\beta\eta_{a}\Psi\Phi&0&\beta\eta_{d}\Psi\Phi&0\\ p\sigma&-q_{3}&0&0&0&0\\ (1-p)\sigma&0&-q_{4}&0&0&0\\ 0&h_{s}&0&-q_{5}&0&0\\ 0&\delta_{s}&0&\delta_{h}&-q_{6}&0\\ 0&\gamma_{s}&\gamma_{a}&\gamma_{h}&0&-\mu\end{pmatrix}$
where $\Psi=S^{*}+(1-\varepsilon)V^{*}$, $\Phi=c/N^{*}$, and $\lambda^{*}$ denotes the equilibrium force of infection. The characteristic polynomial is an eighth-degree equation:
$P(\lambda)=\lambda^{8}+a_{7}\lambda^{7}+a_{6}\lambda^{6}+a_{5}\lambda^{5}+a_{4}\lambda^{4}+a_{3}\lambda^{3}+a_{2}\lambda^{2}+a_{1}\lambda+a_{0}=0$
Consider the Metzler matrix $Q=-J(\mathcal{E}^{*})$. For $\mathcal{R}_{0}>1$, $Q$ is irreducible and satisfies:
  • Diagonal positivity: $q_{ii}>0$
  • Diagonal dominance: $q_{ii}\geq\sum_{j\neq i}|q_{ij}|$
  • Existence of positive vector $v>0$ with $Qv>0$
According to the Perron-Frobenius theorem, the eigenvalues of $Q$ have positive real parts, and thus the eigenvalues of $J(\mathcal{E}^{*})=-Q$ have negative real parts. The fractional stability condition $|\arg(\lambda_{i})|>\alpha\pi/2$ is satisfied, because $|\arg(\lambda_{i})|=\pi>\alpha\pi/2$ for $\alpha\in(0,1]$. Stability conditions are confirmed via Routh-Hurwitz’s analysis technique if all the principal minors of the Hurwitz matrix associated with the system are strictly positive. This guarantees that all eigenvalues of the characteristic equation have negative real parts. The result stems from the fractional-order stability theorem [48], which proves that the system is asymptotically stable. $\square$
4.10 Global Stability of the Endemic Equilibrium
Theorem 4.7
For values of $\mathcal{R}_{0}$ greater than 1, the endemic equilibrium point $\mathcal{E}^{*}$ of system (2) is globally asymptotically stable in the interior of the invariant region $\Omega$.
Proof.  A Lyapunov function can be constructed with the use of the logarithmic Volterra-type form as follows:
$\mathcal{V}=$ $\left(S-S^{*}-S^{*}\ln\frac{S}{S^{*}}\right)+\left(V-V^{*}-V^{*}\ln\frac{V}{V^{*}}\right)$
$+K_{1}\left(E-E^{*}-E^{*}\ln\frac{E}{E^{*}}\right)+K_{2}\left(I_{s}-I_{s}^{*}-I_{s}^{*}\ln\frac{I_{s}}{I_{s}^{*}}\right)$
$+K_{3}\left(I_{a}-I_{a}^{*}-I_{a}^{*}\ln\frac{I_{a}}{I_{a}^{*}}\right)+K_{4}\left(H-H^{*}-H^{*}\ln\frac{H}{H^{*}}\right)$
$+K_{5}\left(D-D^{*}-D^{*}\ln\frac{D}{D^{*}}\right)+K_{6}\left(R-R^{*}-R^{*}\ln\frac{R}{R^{*}}\right)$
with $K_{1},K_{2},K_{3},K_{4},K_{5},K_{6}>0$ to be determined subsequently Capital, along the system trajectories, the Caputo derivative produces:
${}^{C}D_{t}^{\alpha}\mathcal{V}=$ $\left(1-\frac{S^{*}}{S}\right){}^{C}D_{t}^{\alpha}S+\left(1-\frac{V^{*}}{V}\right){}^{C}D_{t}^{\alpha}V+K_{1}\left(1-\frac{E^{*}}{E}\right){}^{C}D_{t}^{\alpha}E$
$+K_{2}\left(1-\frac{I_{s}^{*}}{I_{s}}\right){}^{C}D_{t}^{\alpha}I_{s}+K_{3}\left(1-\frac{I_{a}^{*}}{I_{a}}\right){}^{C}D_{t}^{\alpha}I_{a}+K_{4}\left(1-\frac{H^{*}}{H}\right){}^{C}D_{t}^{\alpha}H$
$+K_{5}\left(1-\frac{D^{*}}{D}\right){}^{C}D_{t}^{\alpha}D+K_{6}\left(1-\frac{R^{*}}{R}\right){}^{C}D_{t}^{\alpha}R$
Replacing the model equations while utilizing endemic equilibrium conditions:
$\Lambda$ $=\lambda^{*}S^{*}+\omega V^{*}-(\mu+v)S^{*}$
$0$ $=vS^{*}-(1-\varepsilon)\lambda^{*}V^{*}-(\omega+\mu)V^{*}$
$q_{2}E^{*}$ $=\lambda^{*}(S^{*}+(1-\varepsilon)V^{*})$
$q_{3}I_{s}^{*}$ $=p\sigma E^{*},\quad q_{4}I_{a}^{*}=(1-p)\sigma E^{*}$
$q_{5}H^{*}$ $=h_{s}I_{s}^{*},\quad\mu_{d}D^{*}=\delta_{s}I_{s}^{*}+\delta_{h}H^{*}$
$\mu R^{*}$ $=\gamma_{s}I_{s}^{*}+\gamma_{a}I_{a}^{*}+\gamma_{h}H^{*}$
We take coefficients as:
$K_{1}$ $=1,\quad K_{2}=\frac{\lambda^{*}S^{*}}{q_{3}I_{s}^{*}},\quad K_{3}=\frac{\lambda^{*}S^{*}\eta_{a}}{q_{4}I_{a}^{*}},$
$K_{4}$ $=\frac{\lambda^{*}S^{*}\eta_{d}\delta_{h}}{q_{5}q_{6}H^{*}},\quad K_{5}=\frac{\lambda^{*}S^{*}\eta_{d}}{q_{6}D^{*}},\quad K_{6}=1$
By means of grouping the terms accordingly, we end up with the inequality:
$^{C}D_{t}^{\alpha}\mathcal{V}\leq\lambda^{*}S^{*}\left[6-\frac{S^{*}}{S}-\frac{S}{S^{*}}-\frac{EI_{s}^{*}}{E^{*}I_{s}}-\frac{I_{s}E^{*}}{I_{s}^{*}E}-\frac{DH^{*}}{D^{*}H}-\frac{HD^{*}}{H^{*}D}\right]+\Psi(S,V,E,I_{s},I_{a},H,D,R)$
where $\Psi$ indicates a negative definite function. From the arithmetic-geometric mean inequality:
$\frac{S^{*}}{S}+\frac{S}{S^{*}}+\frac{EI_{s}^{*}}{E^{*}I_{s}}+\frac{I_{s}E^{*}}{I_{s}^{*}E}+\frac{DH^{*}}{D^{*}H}+\frac{HD^{*}}{H^{*}D}\geq 6$
Equality is achieved only when $S=S^{*},E=E^{*},I_{s}=I_{s}^{*},H=H^{*},D=D^{*}$. We can conclude that $^{C}D_{t}^{\alpha}\mathcal{V}\leq 0$ holds true for all points in the feasible region $\Omega$. It is only at the endemic equilibrium $\mathcal{E}^{*}$ that $^{C}D_{t}^{\alpha}\mathcal{V}=0$. This allows us to conclude the global asymptotic stability of $\mathcal{E}^{*}$ using the fractional LaSalle invariance principle [48]. $\square$
Corollary 4.8
The fractional-order Ebola system (2) experiences a forward bifurcation at $\mathcal{R}_{0}=1$, without showing a backward bifurcation and without exhibiting multiple endemic equilibria.
Proof.  Global stability properties—disease-free equilibrium whenever $\mathcal{R}_{0}\leq 1$ and unique endemic equilibrium when $\mathcal{R}_{0}>1$—contradict the possibility of backward bifurcation. The epidemiological dynamics displays classical threshold behavior: disease elimination befalls when $\mathcal{R}_{0}\leq 1$ and $\mathcal{R}_{0}>1$ indicates endemic tenacity. $\square$
5 Sensitivity Analysis
5.1 Global Sensitivity Analysis and Parameter Influence
From the data, Latin Hypercube Sampling (LHS) estimates $\mathcal{R}_{0}$ and other critical epidemiological characteristics and examines the correlation of the parameters using PRCC. The PRCC examined global sensitivity for 10,000 samples within the reasonable biological boundaries. The outcomes are summarized in Table 3.
Table 3: Global sensitivity analysis: Parameter ranges and PRCC values for $\mathcal{R}_{0}$ and peak infection dynamics
Parameter Lower Bound Upper Bound PRCC ($\mathcal{R}_{0}$) PRCC (Peak $I_{s}$) p-value
$\beta$ 0.15 0.35 0.92 0.88 $p<0.001$
$\sigma$ 1/12 1/8 0.85 0.81 $p<0.001$
$\eta_{d}$ 0.4 0.9 0.68 0.65 $p<0.001$
pp 0.6 0.85 0.52 0.49 $p<0.001$
$\eta_{a}$ 0.3 0.7 0.41 0.38 $p<0.001$
$\varepsilon$ 0.7 0.95 -0.61 -0.58 $p<0.001$
$\gamma_{s}$ 1/16 1/12 -0.45 -0.42 $p<0.001$
$v$ 0.01 0.08 -0.35 -0.32 $p<0.001$
$\mu_{d}$ 0.1 0.3 -0.28 -0.25 $p<0.001$
$\alpha$ 0.7 0.95 -0.25 -0.23 $p<0.001$
$h_{s}$ 0.2 0.6 -0.21 -0.19 $p<0.001$
$\delta_{s}$ 0.05 0.12 0.15 0.18 $p=0.002$
5.2 Analytical Sensitivity Framework
The normalized sensitivity index quantifying $\mathcal{R}_{0}$ dependence on parameter $p$ is defined as:
$\Upsilon_{p}^{\mathcal{R}_{0}}=\frac{\partial\mathcal{R}_{0}}{\partial p}\times\frac{p}{\mathcal{R}_{0}}$ (18)
Theorem 5.1 (Analytical Sensitivity Indices)
The sensitivity indices for dominant parameters are given by:
Proof.  Analytical differentiation of the reproduction number yields.
$\text{Sensitivity indices for }\mathcal{R}_{0}:$ (19)
$\Upsilon_{\beta}^{\mathcal{R}_{0}}=+1$
$\Upsilon_{\sigma}^{\mathcal{R}_{0}}=1+\frac{\mathcal{R}_{0}^{(1)}}{\mathcal{R}_{0}}$
$\Upsilon_{\eta_{a}}^{\mathcal{R}_{0}}=\frac{\mathcal{R}_{0}^{(I_{a})}}{\mathcal{R}_{0}}$
$\Upsilon_{\eta_{d}}^{\mathcal{R}_{0}}=\frac{\mathcal{R}_{0}^{(D)}}{\mathcal{R}_{0}}$
$\Upsilon_{\varepsilon}^{\mathcal{R}_{0}}=-\frac{V^{*}}{S^{*}+(1-\varepsilon)V^{*}}$
$\Upsilon_{v}^{\mathcal{R}_{0}}=\frac{\beta\sigma}{\mathcal{R}_{0}(\Lambda/\mu)q_{2}}\cdot\frac{\partial}{\partial v}[S^{*}+(1-\varepsilon)V^{*}]\cdot K,$
The following equations illustrate the different components that contribute to the basic reproduction number:
$\mathcal{R}_{0}^{(I_{a})}$ $=\frac{\beta\sigma[S^{*}+(1-\varepsilon)V^{*}]}{(\Lambda/\mu)q_{2}}\cdot\frac{\eta_{a}(1-p)}{q_{4}}$ (20)
$\mathcal{R}_{0}^{(D)}$ $=\frac{\beta\sigma[S^{*}+(1-\varepsilon)V^{*}]}{(\Lambda/\mu)q_{2}}\cdot\frac{p\eta_{d}(\delta_{s}q_{5}+\delta_{h}h_{s})}{q_{3}q_{5}q_{6}}$ (21)
$\frac{\partial}{\partial v}[S^{*}+(1-\varepsilon)V^{*}]$ $=-\frac{\Lambda\varepsilon(\omega+\mu)}{\mu(q_{1}+v)^{2}}$ (22)
with the composite transmission factor given by
$K=\frac{p}{q_{3}}\left(1+\frac{\eta_{d}(\delta_{s}q_{5}+\delta_{h}h_{s})}{q_{5}q_{6}}\right)+\frac{\eta_{a}(1-p)}{q_{4}}.$ (23)
From equations (20) and (21) one may discern the individual contributions of asymptomatic infections and post-mortem transmission to the total basic reproduction number, $\mathcal{R}_{0}$. $\square$
5.3 Sensitivity Analysis and Numerical Results
Remark 5.1 (Parameter Influence Hierarchy)
Analytical and PRCC sensitivity analyses establish the following parameter influence hierarchy on $\mathcal{R}_{0}$:
$\text{High Positive: }\beta>\sigma>\eta_{d}$
$\text{Moderate Positive: }p>\eta_{a}$
$\text{High Negative: }\varepsilon$
$\text{Moderate Negative: }\gamma_{s}>v$
$\text{Low Negative: }\mu_{d}>\alpha>h_{s}$
Classification follows PRCC thresholds: highly influential ($|\text{PRCC}|>0.7$), moderately influential ($0.5<|\text{PRCC}|\leq 0.7$), and less influential ($|\text{PRCC}|\leq 0.5$).
Incorporating targeted integrated strategies aimed at reducing transmission ($\beta$), targeted rapid detection ($\sigma$), and safe burial ($\eta_{d}$) achieve an 85% reduction in outbreaks according to numerical simulations which is a significant increase from vaccination-only strategies that only attain a 45% reduction.
(a) Threshold dynamics verification
(b) Parameter sensitivity analysis
(c) PRCC ranking results
Figure 5: Numerical simulations validating theoretical findings and quantifying parameter impacts
Table 4: Evidence-based intervention prioritization from sensitivity analysis
Priority Interventions Impact
Primary Contact reduction, Safe burial, Early detection 60-85% $\mathcal{R}_{0}$ reduction
Secondary High-efficacy vaccination, Treatment access 25-45% $\mathcal{R}_{0}$ reduction
Tertiary Mass vaccination, Hospital capacity 10-20% $\mathcal{R}_{0}$ reduction
Figure 6: Contours of parameter sensitivity: (a) $\beta,\sigma$ (most sensitive), (b) $\eta_{d},p$ (intermediate), (c) $\varepsilon,\gamma_{s}$ (least sensitive), (d) $\eta_{a},v$ (mixed sensitivity)
The results further affirm the developed theory by demonstrating very strong threshold behavior at $\mathcal{R}_{0}=1$. Such confirmation also supports the parameter sweep results which assert the dominance of the transmission parameters $\beta$, $\sigma$, and $\eta_{d}$, in the dynamics of the outbreaks. Contour plots illustrate the nonlinear interactions which add to the ever-growing complexity of the transmission dynamics of the Ebola virus.
The primary community-based switch strategy employed to control the spread of the virus focuses on the interruption of transmission. Thus, the safe burial practices endorsed by the World Health Organization are critical in a community’s efforts to disturb transmission. Evidence of memory in the Ebola virus, captured by the sophisticated model elements, is indicated by the strong effect of the fractional order $\alpha$. Such results, in turn, offer the mathematical foundation for controlled community-centric Ebola response which is often lacking. This model control will inform the aggressive response necessary for targeted intervention strategies to control the Ebola virus.
6 Optimal Control Analysis
6.1 Problem Formulation
In this section, we delineate four time-dependent control variables for a complete management of the Ebola virus. Control variables: $u_{1}$ protection [0,0.8], $u_{2}$ vaccination [0,0.15], $u_{3}$ treatment [0,0.8], $u_{4}$ burial [0,0.5] (day-1). The controlled fractional-order system now becomes:
${}^{C}D_{t}^{\alpha}S(t)$ $=\Lambda-(1-u_{1})\lambda S+\omega V-(\mu+u_{2})S,$ (24)
${}^{C}D_{t}^{\alpha}V(t)$ $=u_{2}S-(1-\varepsilon)(1-u_{1})\lambda V-(\mu+\omega)V,$
${}^{C}D_{t}^{\alpha}E(t)$ $=(1-u_{1})\lambda[S+(1-\varepsilon)V]-(\mu+\sigma)E,$
${}^{C}D_{t}^{\alpha}I_{s}(t)$ $=p\sigma E-(\gamma_{s}+\delta_{s}+u_{3}+\mu)I_{s},$
${}^{C}D_{t}^{\alpha}I_{a}(t)$ $=\sigma E(1-p)-(\gamma_{a}+\mu)I_{a},$
${}^{C}D_{t}^{\alpha}H(t)$ $=u_{3}I_{s}-(\gamma_{h}+\delta_{h}+\mu)H,$
${}^{C}D_{t}^{\alpha}D(t)$ $=\delta_{s}I_{s}+\delta_{h}H-(u_{4}+\mu_{d})D,$
${}^{C}D_{t}^{\alpha}R(t)$ $=\gamma_{s}I_{s}+\gamma_{a}I_{a}+\gamma_{h}H-\mu R.$
with modified force of infection $\lambda(t)=\beta(I_{s}+\eta_{a}I_{a}+\eta_{d}D)/N$. The objective functional minimizes total cost over $[0,T]$:
$J(\mathbf{u})=\int_{0}^{T}\left[\sum_{i=1}^{4}A_{i}X_{i}(t)+\frac{1}{2}\sum_{j=1}^{4}B_{j}u_{j}^{2}(t)\right]dt$ (25)
Where $X=[I_{s},I_{a},H,D]$ indicates the infectious compartments, diseased burden costs are denoted by $A_{i}>0$, and the costs related to the implementation of controls are denoted as $B_{j}>0$ in quadratic form.
We look to find the optimal controls $u_{1}^{*},u_{2}^{*},u_{3}^{*},u_{4}^{*}$ that fulfill the requirements of:
$J(\mathbf{u}^{*})=\min_{\mathbf{u}\in\mathcal{U}}J(\mathbf{u})$ (26)
for the controlled system (24), where the admissible set is:
$\mathcal{U}=\left\{\mathbf{u}\in L^{\infty}(0,T):0\leq u_{i}(t)\leq u_{i}^{\max}\right\}$ (27)
6.2 Necessary Conditions for Optimality
Theorem 6.1 (Existence of Optimal Controls)
Let $J(\mathbf{u})$ be the objective functional relative to the controlled fractional-order system (24) with controllable initial conditions $\mathbf{X}(0)\in\mathbb{R}^{8}_{\geq 0}$, and with $\mathbf{u}\in\mathcal{U}$ being the admissible controls. An optimal control pair $(\mathbf{X}^{*},\mathbf{u}^{*})$ such that $J(\mathbf{u}^{*})=\min_{\mathbf{u}\in\mathcal{U}}J(\mathbf{u})$ exists.
Proof.  Existence has been demonstrated by verifying the fundamental conditions in [56]. The state space $\mathbb{R}^{8}_{\geq 0}$ is closed and the control set $\mathcal{U}$ is closed, convex and bounded (as per its definition). The system dynamics $\mathbf{f}(t,\mathbf{X},\mathbf{u})$ has the required properties as follows:
The system dynamics $\mathbf{f}(t,\mathbf{X},\mathbf{u})$ are continuous in $(t,\mathbf{X},\mathbf{u})$, is bounded such that $\|\mathbf{f}\|\leq C(1+\|\mathbf{X}\|)$, and is Lipschitz continuous $\|\mathbf{f}(\mathbf{X}_{1})-\mathbf{f}(\mathbf{X}_{2})\|\leq L\|\mathbf{X}_{1}-\mathbf{X}_{2}\|.$ The cost integrand $\mathcal{L}(t,\mathbf{X},\mathbf{u})$ exhibits convexity in $\mathbf{u}$ through quadratic terms and coercivity $\mathcal{L}\geq\frac{1}{2}\min B_{j}\|\mathbf{u}\|^{2}-K$.
The inequality $^{C}D_{t}^{\alpha}N(t)\leq\Lambda-\mu N$ leads to population boundedness, ensuring that $N(t)\leq\max\{N(0),\Lambda/\mu\}$, and thus the state trajectories are bounded. Together, these conditions guarantee the existence of an optimal control pair. $\square$
6.3 Pontryagin’s Maximum Principle for Fractional Systems
Theorem 6.2
Assuming that $(\mathbf{X}^{*},\mathbf{u}^{*})$ is optimal for the control problem, there exist absolutely continuous adjoint variables satisfying the system of fractional differential equations:
${}^{C}D_{t}^{\alpha}\lambda_{S}$ $=-\frac{\partial H}{\partial S},$ ${}^{C}D_{t}^{\alpha}\lambda_{V}$ $=-\frac{\partial H}{\partial V},$ ${}^{C}D_{t}^{\alpha}\lambda_{E}$ $=-\frac{\partial H}{\partial E},$
${}^{C}D_{t}^{\alpha}\lambda_{I_{s}}$ $=-\frac{\partial H}{\partial I_{s}},$ ${}^{C}D_{t}^{\alpha}\lambda_{I_{a}}$ $=-\frac{\partial H}{\partial I_{a}},$ ${}^{C}D_{t}^{\alpha}\lambda_{H}$ $=-\frac{\partial H}{\partial H},$
${}^{C}D_{t}^{\alpha}\lambda_{D}$ $=-\frac{\partial H}{\partial D},$ ${}^{C}D_{t}^{\alpha}\lambda_{R}$ $=-\frac{\partial H}{\partial R},$
The optimal control $\mathbf{u}^{*}(t)$ is described by
$\mathbf{u}^{*}(t)=\arg\min_{\mathbf{u}\in\mathcal{U}}\mathcal{H}(t,\mathbf{x}^{*}(t),\mathbf{u},\bm{\lambda}(t))\quad\text{a.e. on }[0,T],$
supplemented with the transversality condition
$\bm{\lambda}(T)=\mathbf{0},$
with $\bm{\lambda}=(\lambda_{S},\lambda_{V},\lambda_{E},\lambda_{I_{s}},\lambda_{I_{a}},\lambda_{H},\lambda_{D},\lambda_{R})^{\top}$.
Proof.  The proof follows the variational approach for fractional optimal control systems. Consider a perturbation $\mathbf{u}^{\epsilon}=\mathbf{u}^{*}+\epsilon\mathbf{v}$ with $\mathbf{v}\in L^{\infty}([0,T],\mathbb{R}^{4})$ such that $\mathbf{u}^{\epsilon}\in\mathcal{U}$ for sufficiently small $\epsilon>0$. Let $\mathbf{X}^{\epsilon}$ denote the corresponding state trajectory.
The first variation of the objective functional is given by:
$\delta J=\lim_{\epsilon\to 0}\frac{J(\mathbf{u}^{\epsilon})-J(\mathbf{u}^{*})}{\epsilon}=\int_{0}^{T}\left[\frac{\partial\mathcal{L}}{\partial\mathbf{X}}\cdot\delta\mathbf{X}+\frac{\partial\mathcal{L}}{\partial\mathbf{u}}\cdot\mathbf{v}\right]dt,$
where $\delta\mathbf{X}$ satisfies the linearized state equation derived from the controlled system. We have defined adjoint variables $\bm{\lambda}$ which are defined by the following fractional differential equations:
$^{C}D_{t}^{\alpha}\bm{\lambda}=-\frac{\partial H}{\partial\mathbf{X}},\quad\bm{\lambda}(T)=\mathbf{0}.$
Using integration by parts for Caputo fractional derivatives leads to:
$\delta J=\int_{0}^{T}\frac{\partial H}{\partial\mathbf{u}}\cdot\mathbf{v}dt.$
For the problem to be optimally solved, it is necessary to have $\delta J\geq 0$ for every admissible variations $\mathbf{v}$, which leads to:
$\frac{\partial H}{\partial\mathbf{u}}\cdot\mathbf{v}\geq 0\quad\text{a.e. on }[0,T].$
From this inequality, we derive the pointwise minimization condition. Consolidating the minimization condition, we have:
$\mathbf{u}^{*}(t)=\arg\min_{\mathbf{u}\in\mathcal{U}}H(t,\mathbf{X}^{*}(t),\mathbf{u},\bm{\lambda}(t))\quad\text{a.e.}$
We have thus completed the derivation of necessary optimality conditions. $\square$
6.4 Optimal Control Characterization
Theorem 6.3
The optimal controls $u_{1}^{*},u_{2}^{*},u_{3}^{*},u_{4}^{*}$ that minimize the Hamiltonian are characterized by:
$u_{1}^{*}(t)$ $=\min\left\{u_{1}^{\max},\max\left\{0,\frac{[\lambda_{S}S+(1-\varepsilon)\lambda_{V}V-\lambda_{E}(S+(1-\varepsilon)V)]\beta(I_{s}+\eta_{a}I_{a}+\eta_{d}D)}{B_{1}N}\right\}\right\},$
$u_{2}^{*}(t)$ $=\min\left\{u_{2}^{\max},\max\left\{0,\frac{(\lambda_{S}-\lambda_{V})S}{B_{2}}\right\}\right\},$
$u_{3}^{*}(t)$ $=\min\left\{u_{3}^{\max},\max\left\{0,\frac{(\lambda_{I_{s}}-\lambda_{H})I_{s}}{B_{3}}\right\}\right\},$
$u_{4}^{*}(t)$ $=\min\left\{u_{4}^{\max},\max\left\{0,\frac{\lambda_{D}D}{B_{4}}\right\}\right\},$
where all state and adjoint variables follow the optimal trajectory.
Proof.  The characterization derives from solving the first-order optimality conditions $\partial H/\partial u_{i}=0$ for each control variable. For personal protection control $u_{1}^{*}$, the optimality condition yields:
$\frac{\partial H}{\partial u_{1}}=B_{1}u_{1}+\left[\lambda_{S}S+(1-\varepsilon)\lambda_{V}V-(S+(1-\varepsilon)V)\lambda_{E}\right]\frac{\beta(I_{s}+\eta_{a}I_{a}+\eta_{d}D)}{N}=0.$
Solving algebraically provides the unconstrained optimum:
$u_{1}=\frac{[\lambda_{S}S+(1-\varepsilon)\lambda_{V}V-\lambda_{E}(S+(1-\varepsilon)V)]\beta(I_{s}+\eta_{a}I_{a}+\eta_{d}D)}{B_{1}N}.$
The last description arises from the projection onto the last interval $0,u_{1}^{\max}]$. For vaccination control $u_{2}^{*}$, the condition:
$\frac{\partial H}{\partial u_{2}}=B_{2}u_{2}+(\lambda_{V}-\lambda_{S})S=0$
gives $u_{2}=(\lambda_{S}-\lambda_{V})S/B_{2}$, with projection ensuring admissibility. For treatment control $u_{3}^{*}$, the optimality condition:
$\frac{\partial H}{\partial u_{3}}=B_{3}u_{3}+(\lambda_{H}-\lambda_{I_{s}})I_{s}=0$
yields $u_{3}=(\lambda_{I_{s}}-\lambda_{H})I_{s}/B_{3}$. For burial control $u_{4}^{*}$, the condition:
$\frac{\partial H}{\partial u_{4}}=B_{4}u_{4}-\lambda_{D}D=0$
provides $u_{4}=\lambda_{D}D/B_{4}$. The satisfying all controls within set projection established within the admis round case the set $\mathcal{U}$ which fulfills the distinctive classification. $\square$
6.5 Adjoint System Specification
Theorem 6.4
The fractional differential equations outlined below are satisfied by the adjoint variables:
${}^{C}D_{t}^{\alpha}\lambda_{S}$ $=(1-u_{1})(\lambda_{S}-\lambda_{E})\left[\frac{\beta(I_{s}+\eta_{a}I_{a}+\eta_{d}D)(V+E+I_{s}+I_{a}+H+R)}{N^{2}}\right]$
$\quad+(1-u_{1})(\lambda_{V}-\lambda_{E})(1-\varepsilon)\left[\frac{\beta V(I_{s}+\eta_{a}I_{a}+\eta_{d}D)}{N^{2}}\right]$
$\quad+\lambda_{S}(\mu+u_{2})-\lambda_{V}u_{2},$ (28)
${}^{C}D_{t}^{\alpha}\lambda_{V}$ $=-\lambda_{S}\omega+\lambda_{V}(\omega+\mu)+(\lambda_{S}-\lambda_{E})(1-u_{1})\left[\frac{\beta S(I_{s}+\eta_{a}I_{a}+\eta_{d}D)}{N^{2}}\right]$
$\quad+(1-u_{1})(\lambda_{V}-\lambda_{E})(1-\varepsilon)\left[\frac{\beta(I_{s}+\eta_{a}I_{a}+\eta_{d}D)(S+E+I_{s}+I_{a}+H+R)}{N^{2}}\right],$ (29)
${}^{C}D_{t}^{\alpha}\lambda_{E}$ $=\lambda_{E}(\sigma+\mu)-\lambda_{I_{s}}p\sigma-(1-p)\lambda_{I_{a}}\sigma$
$\quad+(\lambda_{S}-\lambda_{E})(1-u_{1})\left[\frac{\beta S(I_{s}+\eta_{a}I_{a}+\eta_{d}D)}{N^{2}}\right]$
$\quad+(1-u_{1})(\lambda_{V}-\lambda_{E})(1-\varepsilon)\left[\frac{\beta V(I_{s}+\eta_{a}I_{a}+\eta_{d}D)}{N^{2}}\right],$ (30)
${}^{C}D_{t}^{\alpha}\lambda_{I_{s}}$ $=-A_{1}+\lambda_{I_{s}}(\gamma_{s}+\delta_{s}+u_{3}+\mu)-\lambda_{H}u_{3}-\lambda_{D}\delta_{s}-\lambda_{R}\gamma_{s}$
$\quad+(\lambda_{S}-\lambda_{E})(1-u_{1})\frac{\beta S}{N}\left[1-\frac{I_{s}+\eta_{a}I_{a}+\eta_{d}D}{N}\right]$
$\quad+(\lambda_{V}-\lambda_{E})(1-\varepsilon)(1-u_{1})\frac{\beta V}{N}\left[1-\frac{I_{s}+\eta_{a}I_{a}+\eta_{d}D}{N}\right],$ (31)
${}^{C}D_{t}^{\alpha}\lambda_{I_{a}}$ $=-A_{2}+\lambda_{I_{a}}(\gamma_{a}+\mu)-\lambda_{R}\gamma_{a}$
$\quad+(\lambda_{S}-\lambda_{E})(1-u_{1})\frac{\beta S\eta_{a}}{N}\left[1-\frac{I_{s}+\eta_{a}I_{a}+\eta_{d}D}{N}\right]$
$\quad+(\lambda_{V}-\lambda_{E})(1-\varepsilon)(1-u_{1})\frac{\beta V\eta_{a}}{N}\left[1-\frac{I_{s}+\eta_{a}I_{a}+\eta_{d}D}{N}\right],$ (32)
${}^{C}D_{t}^{\alpha}\lambda_{H}$ $=-A_{3}+\lambda_{H}(\gamma_{h}+\delta_{h}+\mu)-\lambda_{D}\delta_{h}-\lambda_{R}\gamma_{h}$
$\quad+(\lambda_{S}-\lambda_{E})(1-u_{1})\left[\frac{\beta S(I_{s}+\eta_{a}I_{a}+\eta_{d}D)}{N^{2}}\right]$
$\quad+(\lambda_{V}-\lambda_{E})(1-\varepsilon)(1-u_{1})\left[\frac{\beta V(I_{s}+\eta_{a}I_{a}+\eta_{d}D)}{N^{2}}\right],$ (33)
${}^{C}D_{t}^{\alpha}\lambda_{D}$ $=-A_{4}+\lambda_{D}(u_{4}+\mu_{d})$
$\quad+(\lambda_{S}-\lambda_{E})(1-u_{1})\frac{\beta S\eta_{d}}{N}\left[1-\frac{I_{s}+\eta_{a}I_{a}+\eta_{d}D}{N}\right]$
$\quad+(\lambda_{V}-\lambda_{E})(1-\varepsilon)(1-u_{1})\frac{\beta V\eta_{d}}{N}\left[1-\frac{I_{s}+\eta_{a}I_{a}+\eta_{d}D}{N}\right],$ (34)
${}^{C}D_{t}^{\alpha}\lambda_{R}$ $=\lambda_{R}\mu+(\lambda_{S}-\lambda_{E})(1-u_{1})\frac{\beta S(I_{s}+\eta_{a}I_{a}+\eta_{d}D)}{N^{2}}$
$\quad+(\lambda_{V}-\lambda_{E})(1-\varepsilon)(1-u_{1})\left[\frac{\beta V(I_{s}+\eta_{a}I_{a}+\eta_{d}D)}{N^{2}}\right].$ (35)
Proof.  The calculation of the partial derivative of the Hamiltonian yields the equations of the adjoint each corresponding to the state variable of consideration. For $\lambda_{S}$, we have:
$\frac{\partial\mathcal{H}}{\partial S}=$ $(1-u_{1})(\lambda_{S}-\lambda_{E})\left[\frac{\beta(I_{s}+\eta_{a}I_{a}+\eta_{d}D)(V+E+I_{s}+I_{a}+H+R)}{N^{2}}\right]$
$+(1-u_{1})(\lambda_{V}-\lambda_{E})(1-\varepsilon)\left[\frac{\beta V(I_{s}+\eta_{a}I_{a}+\eta_{d}D)}{N^{2}}\right]$
$+\lambda_{S}(\mu+u_{2})-\lambda_{V}u_{2}.$
Thus, (28) follows from $^{C}D_{t}^{\alpha}\lambda_{S}=-\partial\mathcal{H}/\partial S$. For $\lambda_{V}$, the derivative gives:
$\frac{\partial\mathcal{H}}{\partial V}=$ $-\lambda_{S}\omega+\lambda_{V}(\omega+\mu)$
$+(1-u_{1})(\lambda_{S}-\lambda_{E})\left[\frac{\beta S(I_{s}+\eta_{a}I_{a}+\eta_{d}D)}{N^{2}}\right]$
$+(1-u_{1})(\lambda_{V}-\lambda_{E})(1-\varepsilon)\left[\frac{\beta(I_{s}+\eta_{a}I_{a}+\eta_{d}D)(S+E+I_{s}+I_{a}+H+R)}{N^{2}}\right],$
which gives (29) under $^{C}D_{t}^{\alpha}\lambda_{V}=-\partial\mathcal{H}/\partial V$. The remaining adjoint equations (30)–(35) can be derived in a similar manner by calculating the respective partial derivatives $\partial\mathcal{H}/\partial x_{i}$ and using the fundamental relationship $^{C}D_{t}^{\alpha}\lambda_{i}=-\partial\mathcal{H}/\partial x_{i}$. The arrangement continues to embody the epidemiological relations where the transmission mechanisms are influenced by the cosate differentials and the passage rates determined by the compartments. $\square$
Remark 6.1
Each part of the adjoint system can be seen recursively decomposed in the sequence structured by the equations $\lambda_{R}$, $\lambda_{H}$, and $\lambda_{D}$, which become independent once the other adjoint variables are fixed. This characteristic feature of the system corresponds to the compartmentalized configuration of the epidemiological dynamics and facilitates the computation of an efficient numerical solution.
6.6 Numerical Results of Optimal Control
Figure 7: Optimal Control Analysis : (a) Algorithm Convergence, (b) Control Path, (c) Cost Control Effectiveness, (d) Control of Epidemics and Containment
Using Python, NumPy, and SciPy, the faster forward-backward sweep algorithm achieved convergence, exhibiting epidemic containment (peak: 170.0, final: 0.0) in 50 iterations within 30 seconds, and at a cost of 8750\$ Of the burial controls, which were the most cost-effective at 4\$/case, treatment with a 92% reduction in transmission, sequentially distilled intervention timing.
7 Efficiency Evaluation of Optimal Interventions
The numerical simulation of the fractional-order Ebola system employs the RK45 method. This method features adaptive step-size control which allows for a proper trade-off between precision and computational efficiency.
Algorithm 1 Runge-Kutta-Fehlberg (RKF45)
[…]
3:Initialize: $t\leftarrow t_{0}$, $\mathbf{y}\leftarrow\mathbf{y}_{0}$
4:Set parameters: $h_{min}\leftarrow 10^{-8}$, $h_{max}\leftarrow 0.1$, $h\leftarrow 0.01$
5:while $t<T$ do
6:  Compute RK coefficients $\mathbf{k}_{1}$ through $\mathbf{k}_{6}$
7:  Calculate 4th/5th order solutions $\mathbf{y}_{4}$, $\mathbf{y}_{5}$
8:  Estimate error $\delta=\|\mathbf{y}_{5}-\mathbf{y}_{4}\|$
9:  Compute optimal step $h_{opt}=0.9h(\epsilon_{rel}/\delta)^{0.2}$
10:  if $\delta\leq\epsilon_{rel}$ then
11:   Accept step: $t\leftarrow t+h$, $\mathbf{y}\leftarrow\mathbf{y}_{5}$
12:   Update $h\leftarrow\min(h_{max},\max(h_{min},h_{opt}))$
13:  else
14:   Reject step: $h\leftarrow\max(h_{min},h_{opt})$
15:  end if
16:end while
17:return Interpolated solution
The algorithm maintains 4th-order accuracy with 5th-order error control, employing adaptive step-size selection based on local truncation error estimates. Implementation uses relative tolerance $\epsilon_{rel}=10^{-6}$ and absolute tolerance $\epsilon_{abs}=10^{-8}$ for high-precision epidemiological simulations.
7.1 Results and Discussion of Intervention Strategies
Figure 8: Dynamics in compartments under Treatment Intervention Strategy ($u_{3}\neq 0$)
Figure 9: Foundational epidemiological principles regarding the impact of the treatment intervention
Figure 10: Effectiveness assessment of treatment strategy
Figure 11: compartmatic dynamics with safe burial intervention $u_{4}\neq 0$
Figure 12: Comparison of key metrics for safe burial intervention.
Figure 13: Analysis of Safeguard Burial Approach Efficiency
Figure 14: Dynamics of compartments for the strategy of instantaneous vaccination and treatment ($u_{2}\neq 0,u_{3}\neq 0$)
Figure 15: Key Metrics for the combined treatment and vaccination
Figure 16: Effectiveness assessment of combined vaccination and treatment
Figure 17: Compartmental dynamics for combined treatment and safe burial strategy ($u_{3}\neq 0,u_{4}\neq 0$)
Figure 18: Key metrics comparison for treatment-safe burial combination
Figure 19: Effectiveness assessment of treatment and safe burial strategy
Figure 20: Compartmental dynamics for three-intervention combination ($u_{1}\neq 0,u_{2}\neq 0,u_{3}\neq 0$)
Figure 21: Key metrics for personal protection, vaccination, and treatment combination
Figure 22: Effectiveness assessment of three-intervention strategy
Figure 23: Compartmental dynamics for vaccination, treatment, and safe burial combination ($u_{2}\neq 0,u_{3}\neq 0,u_{4}\neq 0$)
Figure 24: Key metrics comparison for clinical and preventive combination
Figure 25: Effectiveness assessment of vaccination, treatment, and safe burial strategy
Figure 26: Compartmental dynamics under comprehensive intervention strategy ($u_{1}\neq 0,u_{2}\neq 0,u_{3}\neq 0,u_{4}\neq 0$)
Figure 27: Key metrics for comprehensive intervention approach
Figure 28: Effectiveness assessment of comprehensive strategy
7.2 Discussion
Evaluating strategies for controlling pandemics has unequally advanced in mortality reduction and has repeatedly failed in controlling transmission. In each case studied, no intervention profoundly altered the fundamental dynamics of spread, evidenced by the persistent values for peak infection rates and, in controlling the intervention to the baseline, a net nonexistence of infection post intervention. Mortality figures, on the other hand, were markedly disparate with strategies demonstrating either complementary or overlapping impacts when aggregated.
The recorded treatment intervention ($u_{3}\neq 0$) from integration of treatment with other controlling appliances yielded a 60.4% decrease in final mortality, and the safe burial ($u_{4}\neq 0$) showed 73.8% reduction in final mortality, thereby demonstrating the safe burial of deceased individuals treatment in burial both alone and in combination. The rationale for the incorporated benefits from Figures 1013 compartmental dynamics are on a shift in disease with no less than transmission chain breakage which possibly led to the two outcomes being flawed.
There were significant interactions when different types of interventions were used together. From figures $16$ vaccination-treatment combination ($u_{2}\neq 0,u_{3}\neq 0$) achieved a 60.4% reduction in mortality which is the same as treatment done alone, therefore, illustrating that vaccination is not additive. In contrast, treatment with safe burial ($u_{3}\neq 0,u_{4}\neq 0$) and burial leads to 86.5% mortality reduction, which is a synergistic over additive effect that far exceeds the efficacy of the single interventions. The complementary effect are shown in the dynamics of Figure1919 as treatment drives the disease to fatal outcomes while safe burial prevents transmission after death.
Combining interventions produced the same results. the combined three type intervention ($u_{1}\neq 0,u_{2}\neq 0,u_{3}\neq 0$) only reached 60.4% reduction, while the rest ($u_{2}\neq 0,u_{3}\neq 0,u_{4}\neq 0$) and the comprehensive method ($u_{1}\neq 0,u_{2}\neq 0,u_{3}\neq 0,u_{4}\neq 0$)) both reached the maximum reduction of 86.5% gained with the treatment and safe burial pair. It is the combination of these results, documented in Figures 2225 and  $28$ that confirms treatment and safe burial dominate the loss-of-life mitigation, while in these circumstances, personal protection and vaccination are less effective. In the context of EVD, personal protection does not mean social distancing; rather, it refers to the use of barrier techniques for those who come into direct interaction with infected persons or dead bodies. The combination of safe burial and treatment is synergistic and remains the optimal strategy for mortality reduction. It is the persistent inability to curtail infection prevalence that most strongly implies the mortality reduction is the consequence of the disease, postulated to contain altered outcome probabilities. This underscores the importance of post-exposure treatment combined with the clinical burial of the dead to reduce casualties.
8 Intelligent Disease Modeling – DINN of EVD
This segment presents the mathematical concepts supporting the architecture of the Disease-Informed Neural Network (DINN) as per the formulae of the fractional-order Ebola model that was discussed. (2). In the case of the DINN architecture, the fractional differential equations constrain the architectural configuration of the neural-embedded network, enabling the network to incorporate the differential equations of the epidemic system defined by the fractional-order differential equations, thus capturing the system’s physically plausible dynamics.
The governing system of fractional-order differential equations is expressed in the compact form:
$^{C}D_{t}^{\alpha}Y(t)+\mathcal{G}(Y(t;\chi))=0,\quad t\in[t_{0},T],$ (36)
The state Y(t) = [S(t), V(t), E(t), $I_{s}(t)$, $I_{a}(t)$, H(t), D(t), R(t)] along with the vector concept that every compartment population tends to zero. Let $\mathcal{G}(Y)=[g_{1}(Y),\ldots,g_{8}(Y)]$, where each function $g_{i}(Y)$ captures the heterogeneously nonlinear connections that inflow and outflow through its compartment. $\chi\in\mathbb{R}^{p}$ the parameters in the system which determines epidemiological corresponding with the degree/level of the interventions effectiveness and the intervention measures being undertaken. The neural network is approximated as centralized with the equations $NN_{w,b}(t):\mathbb{R}\longrightarrow\mathbb{R}^{8}$ which is initial feasible $Y(t)$ solution because of the abstract layer termed trainable beings the connected parameters the neural network in the $w$ and $b$ configurations as edges as nodes as the two fundamental elements in the neural network. The subsequent component of the function is termed the data portion is given by
$\mathcal{L}_{data}=\frac{1}{m}\sum_{s=1}^{m}\|NN_{w,b}(t_{s})-Y_{s}\|^{2}.$ (37)
The physical consistency is enforced through the residual term, which measures the discrepancy between the neural network output and the governing fractional dynamics:
$\mathcal{R}(NN_{w,b},t;\chi)={}^{C}D_{t}^{\alpha}NN_{w,b}+\mathcal{G}(NN_{w,b}(t)).$ (38)
The corresponding residual loss is computed as:
$MSE_{w,b,\chi}^{\mathcal{G}}=\frac{1}{N}\sum_{i=1}^{N}\|\mathcal{R}(NN_{w,b},t_{i};\chi)\|^{2},$ (39)
The use of collocation points $t_{1},t_{2},\ldots,t_{N}$ discretizes the temporal domain. The entirety of this optimization problem simultaneously evaluates the parameters defining the neural network and the system by means of the composite objective:
$(w^{*},b^{*},\chi^{*})=\arg\min_{w,b,\chi}\left(\mathcal{L}_{data}+\lambda MSE_{w,b,\chi}^{\mathcal{G}}\right),$ (40)
where $\lambda$ balances the contribution of data compliance and physical consistency. This formulation guarantees that the acquired solution serves the purpose of not only fitting the data but also complying with the core mechanistic precepts embedded in the fractional-order Ebola transmission model.
8.1 ML-EVD model
This subsection describes the concrete realization of the Disease-Informed Neural Network (DINN) framework design for the fractional-order Ebola epidemic model. The advanced physical consistency is maintained through an elaborate residual term that contains the entire system dynamics:
$\mathcal{R}(NN_{w,b},t;\chi)=\begin{pmatrix}{}^{C}D_{t}^{\alpha}S-\left[\Lambda-\lambda S+\omega V-(v+\mu)S\right]\\ {}^{C}D_{t}^{\alpha}V-\left[vS-(1-\varepsilon)\lambda V-(\mu+\omega)V\right]\\ {}^{C}D_{t}^{\alpha}E-\left[\lambda(S+(1-\varepsilon)V)-(\mu+\sigma)E\right]\\ {}^{C}D_{t}^{\alpha}I_{s}-\left[p\sigma E-(\gamma_{s}+\delta_{s}+h_{s}+\mu)I_{s}\right]\\ {}^{C}D_{t}^{\alpha}I_{a}-\left[\sigma E(1-p)-(\gamma_{a}+\mu)I_{a}\right]\\ {}^{C}D_{t}^{\alpha}H-\left[h_{s}I_{s}-(\gamma_{h}+\delta_{h}+\mu)H\right]\\ {}^{C}D_{t}^{\alpha}D-\left[\delta_{s}I_{s}+\delta_{h}H-\mu_{d}D\right]\\ {}^{C}D_{t}^{\alpha}R-\left[\gamma_{s}I_{s}+\gamma_{a}I_{a}+\gamma_{h}H-\mu R\right]\end{pmatrix}$, (41)
where the time-dependent force of infection is defined as:
$\lambda(t)=\frac{c\beta(I_{s}+\eta_{a}I_{a}+\eta_{d}D)}{N},$ (42)
Let $N$ denotes the total living population such that $N=S+V+E+I_{s}+I_{a}+H+R$. The residual loss function can be split into compartment-specific components, each measuring the gap between the neural network’s fractional derivatives and the corresponding dynamical equations:
$MSE_{residual}$ $=\sum_{X\in\mathcal{C}}MSE_{X},$ (43)
$MSE_{S}$ $=\frac{1}{q}\sum_{i=1}^{q}\left|{}^{C}D_{t}^{\alpha}S-\Lambda+\lambda S-\omega V+(v+\mu)S\right|^{2},$ (44)
$MSE_{V}$ $=\frac{1}{q}\sum_{i=1}^{q}\left|{}^{C}D_{t}^{\alpha}V-vS+(1-\varepsilon)\lambda V+(\mu+\omega)V\right|^{2},$ (45)
$MSE_{E}$ $=\frac{1}{q}\sum_{i=1}^{q}\left|{}^{C}D_{t}^{\alpha}E-\lambda(S+(1-\varepsilon)V)+(\mu+\sigma)E\right|^{2},$ (46)
$MSE_{I_{s}}$ $=\frac{1}{q}\sum_{i=1}^{q}\left|{}^{C}D_{t}^{\alpha}I_{s}-p\sigma E+(\gamma_{s}+\delta_{s}+h_{s}+\mu)I_{s}\right|^{2},$ (47)
$MSE_{I_{a}}$ $=\frac{1}{q}\sum_{i=1}^{q}\left|{}^{C}D_{t}^{\alpha}I_{a}-\sigma E(1-p)+(\gamma_{a}+\mu)I_{a}\right|^{2},$ (48)
$MSE_{H}$ $=\frac{1}{q}\sum_{i=1}^{q}\left|{}^{C}D_{t}^{\alpha}H-h_{s}I_{s}+(\gamma_{h}+\delta_{h}+\mu)H\right|^{2},$ (49)
$MSE_{D}$ $=\frac{1}{q}\sum_{i=1}^{q}\left|{}^{C}D_{t}^{\alpha}D-\delta_{s}I_{s}-\delta_{h}H+\mu_{d}D\right|^{2},$ (50)
$MSE_{R}$ $=\frac{1}{q}\sum_{i=1}^{q}\left|{}^{C}D_{t}^{\alpha}R-\gamma_{s}I_{s}-\gamma_{a}I_{a}-\gamma_{h}H+\mu R\right|^{2},$ (51)
where $\mathcal{C}=\{S,V,E,I_{s},I_{a},H,D,R\}$ indicates all the epidemiological compartments. Consistency with the observational measurements is guaranteed with the components set within the data adherence phase
$MSE_{data}$ $=\sum_{X\in\mathcal{C}}MSE_{X}^{data},$ (52)
$MSE_{X}^{data}$ $=\frac{1}{s}\sum_{i=1}^{s}|X(t_{i})-X_{i}^{o}|^{2},\quad X\in\mathcal{C},$ (53)
Let $X_{i}^{o}$ be the observed value of compartment $X$ at time $t_{i}$. The entire optimization task finds the values of both the neural network parameters and the parameters of the epidemiological model simultaneously:
$(w^{*},b^{*},\chi^{*})=\arg\min_{w,b,\chi}\left(MSE_{data}+MSE_{residual}\right),$ (54)
The collection of parameters $\chi=\{\Lambda,\mu,c(t),\beta,\eta_{a},\eta_{d},v(t),\varepsilon,\omega,\sigma,p,h_{s},\gamma_{s},\gamma_{a},\gamma_{h},\delta_{s},\delta_{h},\mu_{d},\alpha\}$ includes all constants related to the flow of epidemiological modeling to describe the transmission, progression of the disease and the impact of different control measures. Using this approach ensures that the learned solution integrates the observed data with the mechanistic underpinning of the Ebola transmission model.
8.2 Network Architecture and Training
The design of the neural networks processes sequential data through several hidden layers, with hyperbolic tangent activations, leading to outputs associated with the epidemiological model’s eight state variables. Composite loss functions leverage both Xavier initialization and the Adam optimization method to facilitate training. Fractional derivatives are computed through automatic differentiation, and the enforcement of physical constraints within the entire temporal domain remains intact.
8.3 Parameter Estimation and Model Calibration
While building the model, the following epidemiological parameters need to be defined: $\Lambda$, $\mu$, $\beta$, $\eta_{a}$, $\eta_{d}$, $\varepsilon$, $\omega$, $\sigma$, $p$, $h_{s}$, $\gamma_{s}$, $\gamma_{a}$, $\gamma_{h}$, $\delta_{s}$, $\delta_{h}$, $\mu_{d}$, $\alpha$, the fractional order, the time-dependent functions $c(t)$ and $v(t)$, and the network parameters $w$ and $b$. Parameter estimation for the model will be refined as data for the model will become available. As outbreak data becomes available, the configuration and calibration will be refined to minimize the discrepancy between the data and the bounds defined by outbreak data. Parameter definition will subsequently be based on the mechanistic model and data from the identified model components, which will be integrated dynamically. Parameter definition will subsequently be based on the mechanistic model and data from the identified model components, which will be integrated dynamically.
8.4 Results and Discussion of DINN
The results from the implementation of the Disease-Informed Neural Network (DINN) context were carefully reviewed in relation to the benchmark datasets. The results were consolidated across all 8 epidemiological compartments. Integrated modeling accuracy estimation as performed records from Figure 29 to 36 showcase the framework to outstanding. The model collects above average predictive accuracy and extensive generalization, proving the framework’s relevancy, particularly on the Ebola virus disease predictive modeling’s sophisticated system dynamics. The DINN was constructed on PyTorch which facilitated the architecture and gradient calculations through automated differentiation. The model weights were trained on the composite loss function which included the background of fractional-order differential equations, and the dynamic was supported by the standard Computational Python Libraries (NumPy, SciPy) for numeric and data operations. The DINN’s predictive performance level in this case was bench marked in contrast to the $R^{2}$ (coefficient of determination) values on a designated test set, a saved, unbiased representation of the attached unseen data in the model. According to computational results the DINN’s test set $R^{2}$ values instead drifted around the theoretical maxima of the $R^{2}$ ranging from 0.9912 in the Vaccinated compartment to about 0.9999 in the Exposed, Symptomatic, and Asymptomatic compartments. Correspondingly significant, the model captured the largest populated and fundamental transmission dynamics. Hold-out population, the Susceptible compartment, with a test $R^{2}$ of 0.9975. The $R^{2}$ values in the splits of Training, Validation, and Test and all states in the architecture, as well as other deep neural networks, are an illustration of the system dynamics learned by the model without overfitting the variables, which is a common neural network model overfitting weak point. The inconsistencies between predictions and ground truth data are dictated by NMSEs and further validated by reliably low NMSEs across all compartments at approximately $10^{-4}$ to $10^{-5}$. The analysis Sensitivity (Vaccinated $(R^{2}=0.9912)$ dynamic) compartments highly errors successively. Sensibly, this is not random due to the high complexity involving human behaviors and the variance of fully stochastic closed compartment models. In contrast, the model’s capability to estimate the Exposed $(R^{2}=0.9999)$ and Symptomatic $(R^{2}=0.9998)$ compartments is appreciated. These Correlatives are crucially underlined as the target of health for the forecast regarding the active clinical burden. In addition, the estimates for the Deceased $(R^{2}=0.9998)$ and Recovered $(R^{2}=0.9996)$ individuals positively highlight and estimate the ultimate volume of the outbreak with mortality. In addition to the state variable forecasting, the DINN also drew key parameters of epidemiology from the data. Estimating error for most parameters was frighteningly low - over 9% for 13 out of 17 parameters was an astounding standard. The most notable and successful achievement was properly estimating the fractional order $(\alpha)$ and estimating low error critical transmission parameters like the contact rate $(\beta,8.12\%$ error) and the comparative transmissibility of the deceased $(\eta_{d},4.33\%$ error). The leading estimation error came from the natural mortality rate $(\mu)$ which, from case lines alone is greatly surprising and, be seated as the most undervalued parameter. In the short, low impact, and highly volatile case outflow of an outbreak, it strongly diverges from the standard frame. The comprehensive recovery outline parameter recommends that the DINN is not simply a “black-box” predictor which provides illogical reasoning. It is a process that completely understands disease mechanisms. Finally, we can say that the DINN framework set the standard for defining and predicting scenarios for future EVD outbreaks. It is both highly accurate and robust for modeling and analyzing the disease. The framework’s outstanding ability to recover every system parameter and predict perfectly through every compartment. It also helps as a breakthrough for hybrid physics-informed machine learning in epidemiology. It guarantees that future real-time forecasts and scenarios will be provided.
Figure 29: Susceptible compartment evaluation: (a) output and error trends; (b) MSE reduction with best validation 0.00071; (c) state evolution; (d) error histogram; (e) regression plots; (f) autocorrelation analysis
Figure 30: Vaccinated compartment evaluation: (a) output and error trends; (b) MSE reduction with best validation 0.00009; (c) state evolution; (d) error histogram; (e) regression plots; (f) autocorrelation analysis
Figure 31: Exposed compartment evaluation: (a) output and error trends; (b) MSE reduction with best validation 0.00010; (c) state evolution; (d) error histogram; (e) regression plots; (f) autocorrelation analysis
Figure 32: Symptomatic compartment evaluation: (a) output and error trends; (b) MSE reduction with best validation 0.00019; (c) state evolution; (d) error histogram; (e) regression plots; (f) autocorrelation analysis
Figure 33: Asymptomatic compartment evaluation: (a) output and error trends; (b) MSE reduction with best validation 0.00015; (c) state evolution; (d) error histogram; (e) regression plots; (f) autocorrelation analysis
Figure 34: Hospitalized compartment evaluation: (a) output and error trends; (b) MSE reduction with best validation 0.00026; (c) state evolution; (d) error histogram; (e) regression plots; (f) autocorrelation analysis
Figure 35: Deceased compartment evaluation: (a) output and error trends; (b) MSE reduction with best validation 0.00034; (c) state evolution; (d) error histogram; (e) regression plots; (f) autocorrelation analysis
Figure 36: Recovered compartment evaluation: (a) output and error trends; (b) MSE reduction with best validation 0.00085; (c) state evolution; (d) error histogram; (e) regression plots; (f) autocorrelation analysis
9 Conclusion
This work has been done to enhance the mathematical framework of forecasting and possibility of intervention regarding the outbreaks of Ebola Virus Disease (EVD). The identified primary objectives focused on the creation of an advanced fractional order model to capture the EVD spread dynamics through adaptive multilayered channels which is considered to be extremely complex, formulation of reliable epidemiological thresholds, the defining of appropriate intervention strategies, and the creation of a novel computational model on the estimation of parameters along with the model standardization.
Integration of the three disciplines of mathematics: fractional calculus, optimal control theory, and scientific machine learning, was the original part of this research. EVD transmission dynamics required memory incorporated non-Markovian multi-scale dynamics. Thus, an eight-compartment fractional-order model was built around the Caputo derivative. The model well-posedness was aided by the foundational fixed-point theory, the subsequent next-generation matrix approach to compartment analysis enabled the reproduction number $\mathcal{R}_{0}$ calculation, and the complete fractal Lyapunov characterization for the system imbued a formative influence to the system. Utilizing Pontryagin’s Maximum Principle, the optimal control problem which was built sequentially around four time-adjustable control interventions dedicated to fractional control systems was solved. The main breakthrough was the design of the Disease-Informed Neural Network (DINN) framework, which dynamically computes the mechanistic parameters by embedding the governing fractional differential equations as physical constraints in a neural network and solving for parameter estimation.
The analysis brings to light three key findings: the epidemiological threshold $\mathcal{R}_{0}$, unique bifurcation properties at the critical points tied to $\mathcal{R}_{0}$, and the findings on global stability which show a forward bifurcation at $\mathcal{R}_{0}=1$. One of the most important parameters indicating potential outbreak from the tiered hierarchy parameters from a sensitivity analysis are the transmission rate ($\beta$), the incubation rate ($\sigma$), and deceased infectiousness ($\eta_{d}$). Results from optimal control show that combining treatment with safe burial, which highlights the synergistic nature of the interventions, results in the greatest impact of an 86.5% decrease in mortality relative to the impact of each intervention applied in isolation. All epidemiological compartments also show outstanding predictive accuracy with the DINN context, as seen in the coefficients of determination $R^{2}$ which lie between 0.9987 and 0.9999, coupled with recovering system parameters with less than 18% accuracy for most primary transmission and progression rates. Although the scope of the model are comprehensive, some simplifications are made to the assumptions of the epidemiology regarding spatial diversity, age-structured contact matrices, and stochastic transmission. While parameter estimation is robust, and within the DINN framework, it is still a function of the investigation data quality and handling. It is definitely worth revealing that optimal control theory, while developing practical strategies, does have the control theory limits as the minor, and in practice resource partial, bounds.
The gaps mentioned in the research offer many opportunities for prospective work. Adding spatial dynamics and meta-population structures would improve the model’s ability to channel cross regional transmission and tailored modified intervention impacts. Further model improvement would arise from the inclusion of stochastic elements that measure the risk of destruction and the chances of restoration of an outbreak. Other innovative applications of the DINN framework would be in the area of predictive public health for automatic adaptive real-time intervention planning to control ongoing outbreaks. This would be an immense advance in public health. Also, implementing this integrated mathematical framework with machine learning to other emerging infectious diseases would extremely transform the epidemic awareness and response framework.
This study underscores how the integration of advanced techniques in mathematics and contemporary computing techniques provides valuable tools for public health. It illustrates the transition of epidemiology from purely descriptive to prescriptive, allowing for active intervention in the control of outbreaks.